### v1

Abstract
The ability of zero-shot translation emerges when we train a multilingual model with certain translation directions; the model can then directly translate in unseen directions. Alternatively, zero-shot translation can be accomplished by pivoting through a third language (e.g., English). In our work, we observe that both direct and pivot translations are noisy and achieve less satisfactory performance. We propose EBBS, an ensemble method with a novel bi-level beam search algorithm, where each ensemble component explores its own prediction step by step at the lower level but they are synchronized by a “soft voting” mechanism at the upper level. Results on two popular multilingual translation datasets show that EBBS consistently outperforms direct and pivot translations as well as existing ensemble techniques. Further, we can distill the ensemble’s knowledge back to the multilingual model to improve inference efficiency; profoundly, our EBBS-based distillation does not sacrifice, or even improves, the translation quality.
1 Introduction
Machine translation is a widely applicable NLP task that translates a text from a source language to a target language Brown et al. (1990); Bahdanau et al. (2015). The Transformer architecture Vaswani et al. (2017) and pretrained large language models Radford et al. (2019); Raffel et al. (2020); Lewis et al. (2020) have largely improved translation performance, especially in the supervised setting, where a model can learn from large volumes of parallel corpora. However, machine translation remains challenging for low-resource languages, because there are not enough data for large neural networks to learn these languages.
We specifically focus on multilingual translation in the zero-shot setting, where the system is required to translate between unseen language pairs. Since collecting parallel data and training individual models for every translation pair are prohibitively expensive, it is common to build a single multilingual system Johnson et al. (2017); Fan et al. (2021) that can perform translation for all language pairs, most of which are zero-shot translation directions with few exceptions (e.g., English). These models work by prepending a language-indicator token, and zero-shot ability emerges as the model generalizes from trained language pairs to unseen ones (Liu et al., 2021; Wicks and Duh, 2022).
The main drawback of such a multilingual model is that they are noisy in the zero-shot setting due to the lack of supervision, and as a result, they tend to generate less satisfactory outputs (Zhang et al., 2020; Liu et al., 2021). Alternatively, zero-shot translation can be performed by pivoting (Wu and Wang, 2007, 2009), where the model first translates the input into a high-resource language such as English, which is then translated to the target language. However, pivoting requires two translation steps, often leading to an accumulation of errors Babych et al. (2007); Gu et al. (2019).
In this paper, we propose a multi-pivot ensemble approach, where we design a novel ensemble method to aggregate direct translation and the predictions made by pivoting through different languages. Our ensemble approach overcomes the drawback of direct and pivot translations as it can utilize these multiple weak models, yielding high-quality output text despite lacking parallel data.
Building an ensemble for text generation is nuanced as it involves a sequence of word predictions. Word-level ensembles aggregate predictions at each generation step, which is usually achieved by averaging the predicted probabilities Sennrich et al. (2016a); Freitag et al. (2017); Shanbhogue et al. (2023). This may not be ideal for zero-shot translation as the predictions are too noisy, making the averaged probabilities overly smooth. On the other hand, minimum Bayes risk decoding (Bickel and Doksum, 2015, MBR) can be considered a sequence-level voting ensemble, but it is only able to select from weak and noisy candidates given by the direct and pivot translations.
To this end, we propose an ensemble decoding algorithm with bi-level beam search (EBBS). Our EBBS performs two levels of beam search at each generation step: at the lower level, beam search is applied individually to each ensemble component; at the upper level, the ensemble maintains a shared beam by voting and synchronizing the candidates (sub-sequences) in individual beams. Unlike word-level ensembles Freitag et al. (2017); Shanbhogue et al. (2023), EBBS does not average the predicted distributions, encouraging individual predictors to explore their own preferences; unlike sequence-level ensembles Kobayashi (2018); Eikema and Aziz (2020), EBBS does not select from a candidate set, and thus is more flexible since voting is performed throughout the generation process.
We conducted experiments on IWSLT (Cettolo et al., 2017) and Europarl (Koehn, 2005), two popular multilingual datasets for zero-shot machine translation. Results show that EBBS can generate high-quality translations and outperform existing ensemble techniques. In addition, we used EBBS-generated outputs as the distillation signal to further finetune our multilingual translation model, which largely improves the efficiency while retaining the performance or even surpassing EBBS on Europarl.
2 Related Work
Zero-shot machine translation. With monolingual data, it is possible to accomplish zero-shot machine translation by training language-specific autoencoders that share the same latent space (Lample et al., 2018a, b; Mohiuddin and Joty, 2020). In a multilingual setting, zero-shot ability emerges when a single model is trained with multiple translation directions, allowing it to generalize to unseen directions (Dabre et al., 2020; Ranathunga et al., 2023; Johnson et al., 2017; Fan et al., 2021).
A key research direction for multilingual translation is to alleviate task interference among different translation directions (Zaremoodi et al., 2018; Wang et al., 2020). Researchers have developed various parameter-sharing schemes Baziotis et al. (2022); Chronopoulou et al. (2023) and applied optimization-based methods to locate language-specific neurons Wang and Zhang (2022); He et al. (2023); such a model is able to better learn individual translation directions while pooling the knowledge of all languages together.
In our work, we use a standard Transformer model following Johnson et al. (2017) and Liu et al. (2021). Our proposed ensemble algorithm EBBS is compatible with the above methods, as it is agnostic to model architectures.
Ensemble learning. Ensemble methods integrate multiple machine learning systems to form a stronger one (Dong et al., 2020; Yang et al., 2023). Bagging, a classic ensemble technique, works by training multiple models with different portions of data and combining their predictions by averaging or voting (Breiman, 1996; Bühlmann and Yu, 2002). Another popular ensemble approach is boosting, where multiple models are trained sequentially, with each subsequent model focusing on the mistakes of the previous ones (Schapire, 2003; Hastie et al., 2009; Natekin and Knoll, 2013). Unfortunately, bagging and boosting are not compatible with our setting, because we build a self-ensemble with only a single model. Alternatively, stacking combines the outputs by training a meta-model (Wolpert, 1992; Ganaie et al., 2022), but this does not apply in our zero-shot setting either because we do not have groundtruth signal to train the meta-model.
These ensemble techniques nevertheless may be applied to supervised text generation (Freitag et al., 2017; Kobayashi, 2018; Hendy et al., 2021). That being said, they are not ideal as they do not take advantage of the sequential nature of sentences.
Unlike previous work, our EBBS performs bi-level beam search, allowing each component to explore its own predictions at every step but synchronizing all individuals by a “soft voting” mechanism. Our approach is specifically suited to the sequence generation process.
3 Approach
In this section, we first explain our ensemble components in §3.1. We then propose EBBS, a novel ensemble decoding algorithm in §3.2. Finally, we describe in §3.3 knowledge distillation with EBBS-decoded outputs to improve efficiency.
3.1 Ensemble Components
In this work, we focus on zero-shot multilingual machine translation, which requires a system to perform translations for multiple languages, where some translation directions are unseen.
Specifically, our multilingual model is an encoder–decoder Transformer with a byte pair encoding tokenizer (Sennrich et al., 2016b) shared among all languages. The encoder can capture the semantics of tokens in any language, whereas the decoder translates the encoded text into the desired language based on a target-language indicator Johnson et al. (2017); Fan et al. (2021).
We follow the standard English-centric training (Johnson et al., 2017; Liu et al., 2021), where the multilingual model is trained with parallel data with English on one of the sides (e.g., German-to-English and English-to-Romanian). The zero-shot ability emerges during such training, and the model is able to translate directly between unseen language pairs (e.g., German-to-Romanian), known as direct translation Dabre et al. (2020); Ranathunga et al. (2023). An alternative approach is pivot translation, where the multilingual model performs two translations using a high-resource language as a pivot (e.g., first translating German to English, and then English to Romanian).
However, both direct and pivot translations have major weaknesses: the quality of direct translation tends to be low due to the lack of parallel data, whereas pivot translation suffers from an error accumulation as it requires two translation steps (Babych et al., 2007; Gu et al., 2019).
In this paper, we would like to build a self-ensemble of the multilingual model to boost translation quality, where direct and pivot translations are our ensemble components.
However, standard ensemble methods such as averaging and voting may not work well for text generation. Voting, for example, chooses the most voted prediction, but in text generation, the components’ votes often do not share anything in common, because there could be tens of thousands of tokens in the vocabulary. An averaging ensemble, on the other hand, averages the predicted distributions of all components, potentially leading to an overly smooth distribution. As reported in a pilot study in the appendix of Fan et al. (2021), a multilingual averaging ensemble only achieves a small improvement over direct translation. Overall, ensemble methods have not been widely explored for zero-shot multilingual translation in previous literature.
3.2 Our Proposed EBBS Algorithm
Figure 1: Overview of our EBBS algorithm.
Input: $\mathbf{x}$: input sentence;   $Z$: beam size
$K$: number of scorers;   $p_{1},\cdots,p_{K}$: scorers
1 $H\leftarrow\emptyset$triangleright candidate outputs
2 $B_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright upper-level beam
3 for $t=1,2,\cdots$ do
4       $\triangleright$ lower level: individual beam search
5       for $\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1}$ do
6             for $k=1,\cdots,K$ do
7                   $B^{\prime}_{t,k}\leftarrow\emptyset$triangleright lower-level beam
8                   for $\mathrm{y}\in V$ do
9                         $p^{\prime}\leftarrow p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$
10                         $B^{\prime}_{k}.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p\cdot p^{\prime}\rangle)$
11                  $B^{\prime}_{t,k}\leftarrow B^{\prime}_{t,k}.\operatorname{top}(Z)$
12
13      $\triangleright$ upper level: beam synchronization
14       $D\leftarrow\text{empty dictionary}$
15       for $k=1,\cdots,K$ do
16             for $\langle\mathbf{y},p\rangle\in B^{\prime}_{t,k}.\operatorname{top}(Z)$ do
17                   if $\mathbf{y}\in D$ then
18                         $D[\mathbf{y}]\leftarrow p+D[\mathbf{y}]$
19                  else
20                         $D[\mathbf{y}]\leftarrow p$
21
22
23      $B_{t}\leftarrow D.\operatorname{top}(Z)$
24       $\triangleright$ check for termination
25       for $\langle\mathbf{y},p\rangle\in B_{t}$ do
26             if $\mathrm{y}_{t}=\textsc{eos}$ then
27                   $H.\operatorname{add}(\langle\mathbf{y},p\rangle)$
28                   if $|H|=Z$ then
29                         return $H.\operatorname{top}(1)$
30
31
32
Algorithm 1 Our EBBS Algorithm
To this end, we propose an ensemble with bi-level beam search (EBBS), a novel decoding framework that enables different ensemble components to collaborate and vote on each other’s partial generations with two levels of beam search.
At the lower level, each ensemble component performs beam search individually, exploring its own preferred regions of the sentence space. At the upper level, EBBS synchronizes the lower-level beam candidates through a voting mechanism, only keeping the most promising partial generations in a shared, upper-level beam. This allows the ensemble components to vote out spurious partial candidates and improve zero-shot translation performance.
Concretely, we assume there are $K$ ensemble components $p_{1},\cdots,p_{K}$, each predicting the probability of the next word given some prefix.
For the 00th decoding step, EBBS initializes a shared, upper-level beam $B_{0}=\langle\textsc{bos},1\rangle$, suggesting that a sequence is forced to start with a special token bos with probability $1$.
For step $t$, each ensemble component performs a lower-level beam search individually, based on the prefixes in the last step’s shared beam $B_{t-1}$:
$B_{t,k}^{\prime}=$ $\operatorname{top-}\!Z\{\ \langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},\ \ p\cdot p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})\rangle:$
$\hskip 37.5807pt\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1},\ \ \mathrm{y}\in V\}$ (1)
for $k=1,\cdots,K$. Here, $\operatorname{top-}\!Z$ selects $Z$-many sequences with the highest probabilities, $\oplus$ represents string concatenation, $V$ is the vocabulary, and $p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$ is the $k$th ensemble component’s predicted probability at step $t$ given the prefix $\mathbf{y}_{1:t-1}$ and input $\mathbf{x}$.
At the upper level, EBBS synchronizes the individual beams $B_{t,k}^{\prime}$, for $k=1,\cdots,K$, into a single, shared, upper-level beam through a soft voting mechanism, where the candidate set $C_{t}$ is the union of the sequences in individual beams:
$C_{t}=\bigcup\nolimits_{k}\{\mathbf{y}:\langle\mathbf{y},p\rangle\in B_{t,k}^{\prime}\}$ (2)
Then, we evaluate each candidate in $C_{t}$ and compute its overall vote as the sum of the probabilities.
[…]
In this way, the upper level synchronizes every ensemble component with the shared beam $B_{t}$ for the next step of generation.
Intuitively, our voting scheme gives an ensemble component $Z$-many votes, each weighted by the predicted probability. The votes (probabilities) are then tallied (summed) for each candidate to form the upper-level beam. Our bi-level beam search terminates when we have $Z$-many terminated sequences in the shared beam, and returns the sequence with the highest score [^1] For selecting the final output, we follow standard implementations and normalize the joint probabilities by length, i.e., taking the geometric mean of step-wise probabilities (Wolf et al., 2019; Ott et al., 2019). Otherwise, beam search algorithms are often biased towards short sequences. as the ensemble output. We provide the detailed pseudocode for EBBS in Algorithm 1 and an illustration in Figure 1.
Discussion. As shown in Algorithm 2 (Appendix A), traditional beam search keeps a fixed-size beam of high-likelihood partial sequences. It is tempting to combine such vanilla beam search with multiple predictors by directly averaging their probabilities $p(\mathbf{y}|\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}p_{k}(\mathbf{y}|\mathbf{x})$ as the score for beam search (Line 8, Algorithm 2), which has been experimented in previous work Sennrich et al. (2016a); Shanbhogue et al. (2023).
However, our intuition suggests that such an approach may suffer from the over-smoothing problem Wei et al. (2019); Wen et al. (2023b): when multiple translations (known as modes) are plausible given an input, the ensemble process will overly smooth out the modes by probability averaging.
By contrast, EBBS allows each ensemble component to explore its own mode (Lines 4–11, Algorithm 1). In Figure 1, for example, the top sequence yields two plausible next tokens, suggested by each component in the lower level. Their probabilities are not smoothed out in our approach, unlike averaging ensembles. The upper level performs soft voting (Lines 12–19, Algorithm 1) so as to maintain tractable inference.
3.3 EBBS-Based Distillation
To improve inference efficiency, we may perform knowledge distillation based on the outputs of our EBBS algorithm. In particular, we follow Kim and Rush (2016) and apply a sequence-level knowledge distillation loss, treating the output $\hat{\mathbf{y}}$ of our ensemble (called a teacher) as the pseudo-groundtruth for finetuning the multilingual translation model (called a student):
$\mathcal{L}_{\text{KD}}=-\sum_{t=1}^{|\hat{\mathbf{y}}|}\log p(\hat{\mathrm{y}}_{t}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})$ (4)
Our self-distillation is an ensemble-then-distill process. This differs from a straightforward practice of multi-teacher distillation, where the student learns from the union of teachers’ outputs Wu et al. (2021). The commonly applied cross-entropy loss is known to yield overly smooth distributions Wen et al. (2023a, b), and the problem becomes more severe with multiple teachers, leading to unsatisfactory performance of union distillation Shayegh et al. (2024). By contrast, our approach provides the student with a consolidated pseudo-groundtruth translation for an input, causing less confusion during the distillation process especially when teachers disagree with each other.
4 Experiments
4.1 Settings
Datasets. We evaluated EBBS on two popular benchmark datasets for zero-shot machine translation: IWSLT (Cettolo et al., 2017), which contains 4 languages (with English) and 6 zero-shot directions; and Europarl v7 (Koehn, 2005), which contains 9 languages and 56 zero-shot directions.
[…]
Model architecture. We replicated Liu et al. (2021) and trained a multilingual system as our base model. Specifically, we had a 5-layer encoder–decoder Transformer for IWSLT, but used an 8-layer Transformer for Europarl to accommodate more training data and languages. We also followed Liu et al. (2021) and removed residual connections in certain middle layers of the encoder (2nd layer for IWSLT; 2nd and 4th layers for Europarl). Liu et al. (2021) show weakening positional correspondence between layers can significantly improve zero-shot performance, which was also observed in our replication experiments.
More details about our experimental setup can be found in Appendix B.
4.2 Competing Methods
We compare our EBBS against the following methods, including pivoting and different ensembles.
English pivoting. Pivoting is commonly applied for translation (Wu and Wang, 2007, 2009; Vamvas and Sennrich, 2022), and we used English as the pivot in our setting because we have parallel data for translations both from and to English.
Word-level averaging ensemble. Averaging is one of the most widely used ensemble techniques in text generation (Sennrich et al., 2016a; Freitag et al., 2017; Shanbhogue et al., 2023). Essentially, the ensemble components’ probabilities are first averaged before being fed to the standard beam search (Algorithm 2).
Word-level voting ensemble. The voting ensemble, common in classification tasks, picks the output class based on the number of votes from ensemble components (given by $\operatorname{argmax}$). However, voting is not common in text generation, because $\operatorname{argmax}$ may select completely different words by the ensemble components due to the large vocabulary size, making voting ineffective. As a remedy, we pick the word by the highest probability when there is a tie for votes.
Sequence-level voting ensemble. Minimum Bayes risk decoding is originally designed as a single-model decoding algorithm, where it selects a sequence from a set of beam search results based on similarity (Eikema and Aziz, 2020; Müller and Sennrich, 2021). Here, we use it as a sequence-level ensemble technique, where the candidates are the output sequences from different ensemble components. Let $C=\{\mathbf{y}_{1},\cdots,\mathbf{y}_{K}\}$ be the set of candidate outputs given by $K$ ensemble components. The best output is selected as
$\mathbf{y}^{*}=\operatorname*{argmax}\limits_{\mathbf{y}\in C}\sum_{\mathbf{y}^{\prime}\in C\setminus\{\mathbf{y}\}}\text{BLEU}(\mathbf{y},\mathbf{y}^{\prime})$ (5)
where $\text{BLEU}(\mathbf{h},\mathbf{r})$ computes the BLEU score between a hypothesis $\mathbf{h}$ and a reference $\mathbf{r}$. In essence, MBR selects an output that resembles others most, using BLEU as the similarity metric.
4.3 Results and Analysis
IWSLT # Model Average it-nl it-ro nl-it nl-ro ro-it ro-nl
1 Direct translation (Liu et al., ${}^{\dagger}$ 17.7 18.5 17.8 17.9 15.5 19.6 16.8
[…]
3 Pivoting (en) 16.19 17.49 15.09 16.79 13.05 18.34 16.37
4 Word-level averaging ensemble 16.52 16.48 16.49 17.53 13.80 19.07 15.77
[…]
6 Sequence-level voting ensemble (MBR) 16.66 16.54 16.51 17.72 13.64 19.58 15.98
[…]
Europarl # Model Average da-de da-es da-fi da-fr da-it da-nl da-pt
1 Direct translation (Liu et al., ${}^{\dagger}$ 26.9 24.2 33.1 18.1 30.6 26.1 26.3 29.9
[…]
4 Word-level averaging ensemble 27.76 26.21 33.64 18.88 31.09 26.71 27.40 30.37
[…]
7 EBBS (ours) 28.36 26.32 34.28 19.43 31.97 27.67 27.78 31.08
Table 1: Main results on the IWSLT and Europarl datasets. The best results are in bold, and the second best results are underlined. ${}^{\dagger}$ indicates cited results; others were obtained by our experimentation.
Main results. We first replicated the base multilingual model Liu et al. (2021). As shown in Rows 1–2, Table 1, the results are generally close, indicating that the replication is successful and ready for our ensemble research. Further, we tried English pivoting (Row 3), a common zero-shot translation method. In our experiments, we find that it does not outperform direct translation, as pivoting methods may suffer from the error accumulation problem since it requires two steps of translation.
We then compare different ensemble techniques, including the proposed EBBS and alternative ensemble methods. Specifically, we notice that IWSLT contains four languages (with English); thus we have two available pivoting directions (excluding source and target), which, along with direct translation, are our three ensemble components. For Europarl, it contains nine languages; for performance and efficiency concerns (to be shown in Figure 2), we also consider three translation paths as our ensemble components: direction translation, English pivoting, and a second pivot. [^2] We use the first available language in the order of Spanish, German, and French. For example, Spanish-to-German translation will have to use French as the pivot. These languages are chosen because they have the most content on the Internet according to the Web Technology Surveys (https://w3techs.com/technologies/overview/content_language), allowing us to better simulate real-world applications.
We start by studying the standard ensemble technique of word-level averaging (Row 4), which has been used in previous translation research (Freitag et al., 2017). In particular, our averaging ensemble performs worse than direct translation on IWSLT, but is slightly better on Europarl. Our results are not consistent to Freitag et al. (2017), who show a word-level averaging ensemble of random seeds can improve performance. This is understandable because models trained with random seeds exhibit similar behavior, and averaging their probabilities achieves a denoising effect. However, our ensemble components are drastically different in terms of their strengths and expertise, because they learn diverse translation patterns from different languages. Thus, word averaging fails to improve translation quality in our setting.
Alternatively, voting ensembles can also be applied, at either the word level or the sequence level. As seen, word-level voting is not effective, as it does not achieve significant improvements (Row 5). This is expected because the voted words (top predictions) by the ensemble components may not overlap due to the large vocabulary size. In such cases, the algorithm defaults to choosing the word with the highest probability, causing the ensemble to follow the most peaked distributions.
Sequence-level voting should also be done in a soft manner, and minimum Bayes risk (MBR) decoding can be thought of as using a Bayes risk to softly “vote” the candidate outputs. As seen from Row 6, such a method works relatively well on Europarl, achieving the second-highest performance across all competing methods; however, it works poorly on the IWSLT dataset. The main drawback of sequence-level voting is that it can only select one of the ensemble components’ outputs. This may not work well when the individual ensemble components are weak, as in our zero-shot setting, and especially with the IWSLT dataset due to a relatively small training set. Such a selective sequence-level ensemble cannot integrate different expertise of its components during generation.
Dataset Method Avg. BLEU Wins Losses
IWSLT Direct 17.28 2 4
Ensemble 18.23 4 2
Europarl Direct 27.85 4 52
Ensemble 28.44 52 4
Overall Direct 26.83 6 56
Ensemble 27.45 56 6
$p$-value 3e-11
Table 2: Performance statistics on all 66 zero-shot directions in both datasets. $p$-value is given by a two-sided binomial test.
Unlike existing ensemble methods, our EBBS algorithm achieves the highest performance in most directions on both datasets. Notice that Europarl contains 56 zero-shot directions; due to the space limit, we can only present in Table 1 the first seven directions based on the order provided by the dataset. Table 2 further shows a pairwise comparison against direct translation (a strong baseline in our experiment) in all zero-shot directions. As seen, EBBS achieves higher performance in 56 out of 62 cases across two datasets, showing strong statistical evidence for its effectiveness, with a $p$-value of 3e-11 in a two-sided binomial test.
Voting scheme BLEU BLEU1 BLEU2 BLEU3 BLEU4
None (beam search) 25.33 56.32 30.08 19.01 12.78
Total-sum 25.27 56.67 30.27 19.07 12.65
[…]
Top-$Z$ sum (ours) 26.10 57.07 31.00 19.76 13.28
Table 3: Comparison of different ensemble variants. We used Italian-to-Dutch translation in the Europarl dataset as a testbed due to the limit of time and resources.
Analysis of voting methods in EBBS. In our EBBS algorithm, the lower-level beams are synchronized into a shared upper-level beam by voting. Specifically, EBBS uses a sum-voting mechanism, where we add the ensemble components’ probabilities for each appearance of a candidate in the lower-level beam, shown in Eqn. (3). Here, we analyze a few alternative voting methods for EBBS.
If EBBS adopts total-sum voting, it still uses lower-level beams to find candidates, but adds all components’ probabilities together. This is equivalent to applying the common averaging ensemble to the top-$Z$ candidates. However, it differs from our approach, because in total-sum voting, a component will vote even if the candidate does not appear in its own lower-level beam; the probability after voting in Eqn. (3) is substituted with $\frac{1}{K}\sum_{k}p_{k}(\mathbf{y}|\mathbf{x})$. As shown in Table 3, EBBS with total-sum voting performs worse than direct translation, suggesting the importance of ignoring the components whose lower-level beam does not contain the candidate. This is analogous to nucleus sampling Holtzman et al. (2019), where the long tail of a distribution is mainly noise and should be ignored.
Other voting schemes that EBBS may use include 0/1 voting and max voting. The former selects the candidates that appear most in the lower-level beams, disregarding the probability values (unless for ties); the latter chooses the maximum probability across the lower-level beams, which gives preference to sequences through a maximization bias (Hasselt, 2010; van Hasselt et al., 2016). As seen in Table 3, EBBS performs relatively well with both of these voting schemes, achieving a decent improvement over the baseline approach; however, their performance is consistently worse than our sum voting scheme.
Overall, our bi-level beam search ensemble is effective with different voting schemes (except for the total-sum voting), and our sum voting works the best among these variants.
Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
Direct 25.33 56.32 30.08 19.01 12.78
Pivoting (en) 25.08 56.76 30.29 19.06 12.66
Pivoting (es) 24.40 55.38 29.08 18.22 12.09
Pivoting (pt) 24.34 55.46 29.02 18.13 12.02
Pivoting (fr) 24.20 55.41 29.02 18.00 11.84
Pivoting (de) 23.65 55.33 28.69 17.67 11.54
Pivoting (da) 23.12 54.81 27.96 17.12 11.12
Pivoting (fi) 20.74 53.54 26.10 15.43 9.79
[…]
Top-3 ensemble 26.10 57.07 31.00 19.76 13.28
Table 4: The performance of individuals and their ensembles for Italian-to-Dutch translation on Europarl.
[…]
Notably, EBBS can achieve a solid improvement (~0.6 BLEU points) over the baseline despite including weak ensemble components (e.g., Finnish pivoting, which is almost 5 BLEU points lower than direct translation). If we limit the ensemble components to the strong ones (top-3), EBBS can achieve an even higher improvement of ~0.8 BLEU points. This analysis highlights the robustness of our EBBS algorithm.
Figure 2: Analysis of the number of ensemble components for Italian-to-Dutch translation on Europarl.
We further study how EBBS performs with different numbers of ensemble components. Specifically, we analyze two incremental ensemble settings: best-to-worst and worst-to-best. In both cases, we start with direct translation; then we incrementally add the next “best” or “worst” pivot translation according to Table 4.
Figure 2 shows the trends of incremental ensembles. If we add best pivot directions, the performance peaks at three ensemble components; interestingly, the inclusion of weaker components does not affect EBBS much. On the other hand, adding the worst pivot translation leads to an immediate drop of 1.6 BLEU points, which then largely recovers with the second pivot. This is reasonable because the worst pivot (Finnish) is 4.6 BLEU points lower than direct translation, and EBBS cannot decide on which of the two ensemble components to trust; despite this, the performance of EBBS is still much better than the average performance of the components. With a second pivot, it gives a third “opinion” when the first two components “disagree.” The performance continues to rise if more and stronger components are added. In fact, our ensemble even surpasses the baseline with 4 weakest pivot translations, each of which is at least 1 BLEU point lower than the baseline. This demonstrates that EBBS is flexible and works well with both strong and weak ensemble components.
Distillation results. We apply distillation techniques to our ensemble outputs to speed up inference. We follow the same ensemble setting (three components) as in Table 1, but only focus on Italian-to-Dutch translation because generating the training samples for all 66 zero-shot directions (each containing more than 100K samples) requires significant computing resources.
We compare our EBBS-based distillation with two alternatives. A straightforward way to distill from multiple teachers (here, translation paths) is to take the union of teachers’ outputs for training. We also include self-distillation, which does not utilize different translation paths but generates samples by standard beam search for self-training.
As seen in Table 5, self-distillation improves zero-shot translation on both IWSLT and Europarl. This is because the base model has never learned from parallel data between Italian and Dutch, and the self-distillation process enables the multilingual model to further refine its translation capabilities in this zero-shot direction. Similarly, we observe that union distillation also improves performance to some degree, but it is not better than self-distillation. This is consistent with the challenges of one-to-many learning seen in previous work (Wen et al., 2023a, b). In our case, learning from the outputs of multiple teachers in a straightforward manner may confuse the student, failing to consolidate and transfer the teachers’ knowledge.
In comparison, EBBS-based distillation consistently outperforms both self-distillation and union distillation, achieving near-EBBS performance on Europarl. For IWSLT, our model even outperforms EBBS, as IWSLT is a small dataset and the high-quality outputs generated by EBBS can largely improve the base model (no distillation) by 2.7 points.
Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
IWSLT EBBS 19.52 51.87 25.12 13.88 8.02
[…]
Union distillation 17.80 49.21 23.01 12.51 7.10
EBBS distillation 20.13 53.20 26.06 14.33 8.26
Europarl EBBS 26.10 57.07 31.00 19.76 13.28
Direct No distillation 25.33 56.32 30.08 19.01 12.78
Self-distillation 25.44 56.54 30.28 19.13 12.79
[…]
EBBS distillation 25.92 56.76 30.68 19.57 13.24
Table 5: Comparison of various distillation methods for Italian-to-Dutch translation.
We provide an efficiency analysis in Appendix D, showing that the inference time of EBBS is linear to the number of ensemble components. Overall, our EBBS-based distillation achieves a significant speed up while maintaining, or even improving, translation quality.
Additional results. We show additional results in the appendices. C: Analysis of beam size, D: Analysis of inference efficiency, E: Entropy of distilled models, and F: Case study.
5 Conclusion
In this work, we address ensemble-based zero-shot machine translation by directly translating and pivoting through different languages. We further design a novel bi-level beam search algorithm (called EBBS) for decoding. We evaluated EBBS on two popular zero-shot translation datasets, IWSLT and Europarl. Results show that EBBS outperforms existing ensemble techniques, and that the high-quality translations produced by EBBS can be used for distillation to improve translation efficiency (and sometimes also quality over EBBS).
6 Limitations
Our work features both algorithmic design and empirical effectiveness, but may also have limitations.
First, our work focuses on zero-shot translation, due to the background of our project. Nevertheless, the EBBS algorithm that we have developed is a general decoding method and is potentially applicable to various text generation tasks. We are happy to explore this direction as future work, such as building ensembles of large language models.
Second, our distillation experiments were conducted on Italian-to-Dutch translation because it appears in both IWSLT and Europarl datasets. We are unable to perform all-direction distillation in this project, because Europarl, say, has 56 zero-shot directions, and distillation techniques require extensive computing resources to generate pseudo-groundtruth sentences. Nevertheless, our distillation results are consistent on both datasets, and therefore, not including all-direction distillation does not hurt the validity of our method. Moreover, our work points to interesting future work, such as iteratively applying our ensemble-then-distill procedure to build a self-improving multilingual machine translation system.
Third, we have mainly used BLEU (in particular, SacreBLEU) scores to evaluate translation quality, which is standard in multilingual translation research Johnson et al. (2017); Liu et al. (2021); Fan et al. (2021). We are aware of language model-based scores, such as BERTScore Zhang et al. (2019), for evaluating the translation in certain languages, but they may not be appropriate for low-resource languages and are not commonly used in our setting. We have not performed human evaluation either, due to both practical and ethical concerns. Practically, it is difficult to find qualified annotators because our multilingual setting requires an annotator to know a large number of languages in addition to English. Ethically, it is inappropriate to ask annotators to study these languages before or during annotation, which is exhausting and may last for extended periods of time. That being said, our work uses the most standard metrics and results are consistent in most BLEU variants on both datasets, with a strong statistical confidence level based on 62 translation directions. In addition, we have provided case studies to further illustrate how our EBBS works. Therefore, we deem our evaluation ethical, appropriate, and adequate.
Acknowledgment
The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), a Mitacs Accelerate project, the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, and the Digital Research Alliance of Canada (alliancecan.ca). In the title, we used the open-domain artwork “The Great Wave off Kanagawa” by Katsushika Hokusai.
Appendix A Beam Search
We show the standard beam search in Algorithm 2 for a comparison with our proposed EBBS. In general, beam search takes a scorer $p$ as the input and approximately finds the highest-scored sequence, by expanding its search tree with all the vocabulary (Lines 6–9) but only keeping the top-$Z$ partial candidates (Line 10) at each generation step. Unlike EBBS, beam search is not specifically designed to work with multiple scorers, and we show in our main analysis that applying beam search with averaged probabilities of the ensemble components is not an ideal approach for ensemble decoding.
Input: $\mathbf{x}$: input sentence
       $Z$: beam size
     $p$: scorer
1 $H\leftarrow\emptyset$triangleright candidate outputs
2 $B_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright beam candidates
3 for $t=1,2,\cdots$ do
4       $\triangleright$ core of beam search
5       $B\leftarrow\emptyset$
6       for $\langle\mathbf{y}_{1:t-1},p^{\prime}\rangle\in B_{t-1}$ do
7             for $\mathrm{y}\in V$ do
8                   $p^{\prime}\leftarrow p^{\prime}\cdot p(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$
9                   $B.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p^{\prime}\rangle)$
10
11      $B_{t}\leftarrow B.\operatorname{top}(Z)$
12       $\triangleright$ check for termination
13       for $\langle\mathbf{y}_{1:t},p^{\prime}\rangle\in B_{t}$ do
14             if $\mathrm{y}_{t}=\textsc{eos}$ then
15                   $H.\operatorname{add}(\langle\mathbf{y},p^{\prime}\rangle)$
16                   if $|H|=Z$ then
17                         return $H.\operatorname{top}(1)$
18
19
20
Algorithm 2 Beam Search
Appendix B Details of Our Experiments
Dataset details. We evaluated our methods using IWSLT 2017 Cettolo et al. (2017) and Europarl v7 Koehn (2005). Table 6 provides a summary of the languages.
The IWSLT 2017 translation dataset features multilingual data derived from TED talks. We followed previous work and used a standard split for zero-shot evaluation (Dabre and Kurohashi, 2017; Liu et al., 2021). In particular, IWSLT contains English-centric training data for Italian, Dutch, and Romanian, while evaluation is performed in six zero-shot directions. IWSLT is a relatively small dataset, which tests our method’s ability to generalize from few languages.
Code Language IWSLT Europarl
da Danish $\checkmark$
de German $\checkmark$
en English $\checkmark$ $\checkmark$
es Spanish $\checkmark$
fi Finnish $\checkmark$
fr French $\checkmark$
it Italian $\checkmark$ $\checkmark$
nl Dutch $\checkmark$ $\checkmark$
pt Portugese $\checkmark$
ro Romanian $\checkmark$
Table 6: Summary of the language codes.
Europarl is a multilingual dataset crawled from the proceedings of the European Parliament. We again followed previous work (Liu et al., 2021) and evaluated our methods with a standard split for the zero-shot setting, containing English-centric data for eight languages with a total of 56 zero-shot evaluation directions. We specifically adopted the non-overlapping setting: in the original corpus, a sentence may be translated into multiple languages, and the non-overlapping setup chooses only one target translation for each input. This prevents potential data leaking problems. Europarl contains more data and languages than IWSLT, which further tests our method’s ability to generalize across multiple languages.
[…]
In our presentation of beam search and the proposed EBBS, we describe the scorer as the multiplication of step-wise probabilities. In implementation, we used the sum of log-probabilities for numerical stability. Moreover, our EBBS is built on top of the popular fairseq framework (Ott et al., 2019), using their beam search implementation as the backbone. Consequently, we inherit standard beam search implementation techniques such as length normalization and max length constraints, which are not detailed in our pseudocode.
Beam Size BLEU BLEU1 BLEU2 BLEU3 BLEU4
1 24.98 56.75 30.08 18.84 12.46
3 25.99 57.04 30.92 19.65 13.17
[…]
Table 7: Beam size analysis for Italian-to-Dutch translation on the Europarl dataset.
Appendix C Analysis of Beam Size
We analyze the effect of the beam size and show the results in Table 7. As seen, performance rises with a larger beam size and eventually plateaus at around five, which is consistent with the practice of standard beam search Meister et al. (2020). This analysis also shows the robustness of our EBBS algorithm. Based on this experiment and efficiency considerations, we used a beam size of five in our main experiments.
It should be mentioned that in principle our lower-level and upper-level beams need not have the same size. In our experiments, we simply set them to be the same. We have not tried full-grid hyperparameter tuning because the focus of our paper is algorithmic design, and as shown, our proposed EBBS is mostly robust to the beam size.
Figure 3: Inference time analysis on the test set of Italian to Dutch translation from the Europarl dataset. Experiments were conducted on an AMD EPYC 7313 CPU and an NVIDIA RTX A6000 GPU, with an inference batch size of 300 samples.
Appendix D Analysis of Inference Efficiency
We analyze the efficiency of our ensemble approach. As seen in Figure 3, the inference scales almost linearly, which is reasonable as we need to perform inference for all the components. The trend shows that it is computationally feasible to build an ensemble of even more components.
Further, the analysis suggests that our EBB-distilled models achieves a speedup of multiple times depending on the number of ensemble components. Recall from §4.3 that our our EBBS-based distillation is able to transfer the ensemble’s knowledge to a single multilingual model, without hurting, if not improving, the translation quality.
Dataset Method BLEU Entropy
[…]
Direct translation No distillation 17.46 2.46
Self-distillation 18.10 1.62
Union distillation 17.80 1.80
EBBS distillation 20.13 1.70
Europarl EBBS 26.10 -
Direct translation No distillation 25.33 2.06
Self-distillation 25.44 1.44
Union distillation 25.53 1.59
[…]
Table 8: Entropy of various distillation techniques on Italian-to-Dutch translation.
IWSLT
Input
ho sempre creduto che trasformare la paura in divertimento sia il dono della creatività.
(I have always believed that turning fear into fun is the gift of creativity.)
Reference
Ik heb altijd geloofd dat het omzetten van angst in plezier de gift is van creativiteit.
(I have always believed that turning fear into joy is the gift of creativity.)
Direct translation
Ik geloofde altijd dat het transformeren van angst in ontspanning de gift van creativiteit is
(I always believed that transforming anxiety into relaxation is the gift of creativity.)
English-pivoting
Omdat ik altijd geloofde om angst in plezier te transformeren, is het geschenk van creativiteit.
(Because I always believed to transform fear into pleasure is the gift of creativity.)
Romanian-pivoting
In feite hebben we altijd gedacht dat het transformeren van angst in divergentie de gift van creativiteit is.
(In fact, we have always thought that transforming fear into divergence is the gift of creativity)
EBBS
Ik geloofde altijd dat het transformeren van angst in plezier de gift van creativiteit is.
(I always believed that transforming fear into pleasure is the gift of creativity.)
Europarl
Input
si poteva avvertire una forte tensione.
(a strong tension could be felt.)
Reference
Er was veel spanning zichtbaar.
( There was a lot of tension visible.)
Direct translation
Er was grote spanning te ontgaan.
(There was great tension to be escaped.)
English-pivoting
Er zou veel spanningen kunnen zijn ontstaan.
(A lot of tensions could have arisen.)
Spanish-pivoting
Mocht een sterke spanning kunnen worden aangekondigd.
(Should a strong tension can be announced.)
EBBS
Er was veel spanning geweest.
(There had been a lot of tension.)
Table 9: Case studies, where the source language is Italian and the target is Dutch. We provide English interpretations in (italic) for non-English text using Google Translate.
Appendix E Entropy of Distilled Models
We would like to understand why EBBS-based distillation largely outperforms other methods, such as union distillation (§4.3). The hypothesis is that cross-entropy distillation loss with diverse samples may lead to an overly smooth distribution, which in turn would affect the model performance Wen et al. (2023b); Shayegh et al. (2024).
We show the average prediction entropy of our distilled models in Table 8. For some input $\mathbf{x}$ and generation step $t$, the prediction entropy is
$H=-\sum_{\mathrm{y}\in V}$ $\;p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})\log p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})$
Here, a large entropy generally indicates that the model is less certain, producing a more uniform prediction, whereas a low entropy indicates that the model is confident, producing a more peaked distribution.
As seen in Table 8, the model without distillation yields the highest entropy, suggesting that it is uncertain about zero-shot translation probably due to a lack of training signals.
Union distillation trains the model from the union of ensemble components’ outputs. It reduces prediction entropy compared with no distillation, but due to the nature of cross-entropy loss, it remains the highest among all distillation variants. Self-distillation is based on direction translation only, reinforcing the model’s current belief and thus producing the lowest entropy. On the contrary, our EBBS-based distillation achieves a moderate entropy on both datasets.
It should be emphasized that the entropy analysis only merely light on how different distillation methods behave, but the entropy itself does not indicate the quality of a model. We quote BLEU scores from Table 5, which has suggested that our EBBS-based distillation achieves similar or higher performance compared with EBBS, consistently outperforming other distillation methods.
Appendix F Case Study
Table 9 shows examples of direct, pivot, and EBBS translations. As seen, pivot and direction translations are prone to low-quality output, but EBBS enables them to correct each other’s mistakes. In the first example, say, our EBBS generally follows the sentence structure of direct translation, where the Italian word “divertimento” (fun) is mistranslated to the Dutch word “ontspanning” (relaxation), but our EBBS corrects it to “plezier” (pleasure), advocated by English pivoting and voted by all ensemble components.

### v2

@@
-
+Yuqiao Wen1,, Behzad Shayegh1,∗, Chenyang Huang1, Yanshuai Cao2, Lili Mou1,3
 Abstract
@@
 In our work, we observe that both direct and pivot translations are noisy and achieve less satisfactory performance.
-We propose EBBS, an ensemble method with a novel bi-level beam search algorithm, where each ensemble component explores its own prediction step by step at the lower level but they are synchronized by a “soft voting” mechanism at the upper level.
-Results on two popular multilingual translation datasets show that EBBS consistently outperforms direct and pivot translations as well as existing ensemble techniques.
-Further, we can distill the ensemble’s knowledge back to the multilingual model to improve inference efficiency; profoundly, our EBBS-based distillation does not sacrifice, or even improves, the translation quality.
+We propose EBBS, an ensemble method with a novel bi-level beam search algorithm, where each ensemble component explores its own prediction step by step at the lower level but all components are synchronized by a “soft voting” mechanism at the upper level.
+Results on two popular multilingual translation datasets show that EBBS consistently outperforms direct and pivot translations, as well as existing ensemble techniques.
+Further, we can distill the ensemble’s knowledge back to the multilingual model to improve inference efficiency; profoundly, our EBBS-distilled model can even outperform EBBS as it learns from the ensemble knowledge.
+GitHubhttps://github.com/MANGA-UOFA/EBBS
 1 Introduction
-Machine translation is a widely applicable NLP task that translates a text from a source language to a target language Brown et al.
-(1990); Bahdanau et al.
-(2015).
-The Transformer architecture Vaswani et al.
-(2017) and pretrained large language models Radford et al.
-(2019); Raffel et al.
-(2020); Lewis et al.
-(2020) have largely improved translation performance, especially in the supervised setting, where a model can learn from large volumes of parallel corpora.
-However, machine translation remains challenging for low-resource languages, because there are not enough data for large neural networks to learn these languages.
+Machine translation is a widely applicable NLP task that aims to translate a text from a source language to a target language (Brown et al.
+1990; Bahdanau, Cho, and Bengio 2015).
+The Transformer architecture (Vaswani et al.
+2017) and pretrained large language models (Radford et al.
+2019; Lewis et al.
+2020) have largely improved translation performance, especially in the supervised setting (Raffel et al.
+2020), where a model can learn from large volumes of parallel corpora.
+However, machine translation remains challenging for low-resource languages, because there are not enough data for large neural networks to learn these languages (Radford et al.
+2019; Muennighoff et al.
+2023).
 We specifically focus on multilingual translation in the zero-shot setting, where the system is required to translate between unseen language pairs.
-Since collecting parallel data and training individual models for every translation pair are prohibitively expensive, it is common to build a single multilingual system Johnson et al.
-(2017); Fan et al.
-(2021) that can perform translation for all language pairs, most of which are zero-shot translation directions with few exceptions (e.g., English).
-These models work by prepending a language-indicator token, and zero-shot ability emerges as the model generalizes from trained language pairs to unseen ones (Liu et al., 2021; Wicks and Duh, 2022).
-The main drawback of such a multilingual model is that they are noisy in the zero-shot setting due to the lack of supervision, and as a result, they tend to generate less satisfactory outputs (Zhang et al., 2020; Liu et al., 2021).
-Alternatively, zero-shot translation can be performed by pivoting (Wu and Wang, 2007, 2009), where the model first translates the input into a high-resource language such as English, which is then translated to the target language.
-However, pivoting requires two translation steps, often leading to an accumulation of errors Babych et al.
-(2007); Gu et al.
-(2019).
-In this paper, we propose a multi-pivot ensemble approach, where we design a novel ensemble method to aggregate direct translation and the predictions made by pivoting through different languages.
-Our ensemble approach overcomes the drawback of direct and pivot translations as it can utilize these multiple weak models, yielding high-quality output text despite lacking parallel data.
+Since collecting parallel data and training individual models for every translation pair are prohibitively expensive, it is common to build a single multilingual system (Johnson et al.
+2017; Fan et al.
+2021) that can perform translation for all language pairs, most of which are zero-shot translation directions that do not involve a high-resource language (e.g., English).
+These models work by prepending a language-indicator token; the zero-shot translation ability emerges as the model generalizes from trained language pairs and is able to perform direct translation for unseen ones (Liu et al.
+2021; Wicks and Duh 2022).
+The main drawback of such multilingual models is that they are noisy in the zero-shot setting due to the lack of supervision, and as a result, they tend to generate low-quality translations (Zhang et al.
+2020; Liu et al.
+2021).
+Alternatively, zero-shot translation can be performed by pivoting (Wu and Wang 2007, 2009), where the model first translates the input into a high-resource language such as English, which is then translated to the target language.
+However, pivoting requires two translation steps, often leading to an accumulation of errors (Babych, Hartley, and Sharoff 2007; Gu et al.
+2019).
+In this paper, we propose an ensemble approach that aggregates direct and pivot translations in order to build a stronger multilingual translation model from weak ones.
 Building an ensemble for text generation is nuanced as it involves a sequence of word predictions.
-Word-level ensembles aggregate predictions at each generation step, which is usually achieved by averaging the predicted probabilities Sennrich et al.
-(2016a); Freitag et al.
-(2017); Shanbhogue et al.
-(2023).
+Word-level ensembles aggregate predictions at each generation step, which is usually achieved by averaging the predicted probabilities (Sennrich, Haddow, and Birch 2016a; Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al.
+2023).
 This may not be ideal for zero-shot translation as the predictions are too noisy, making the averaged probabilities overly smooth.
-On the other hand, minimum Bayes risk decoding (Bickel and Doksum, 2015, MBR) can be considered a sequence-level voting ensemble, but it is only able to select from weak and noisy candidates given by the direct and pivot translations.
+On the other hand, minimum Bayes risk decoding (MBR) (Bickel and Doksum 2015) can be considered a sequence-level voting ensemble, but existing MBR methods are only able to select from weak and noisy candidates given by the direct and pivot translations.
 To this end, we propose an ensemble decoding algorithm with bi-level beam search (EBBS).
-Our EBBS performs two levels of beam search at each generation step: at the lower level, beam search is applied individually to each ensemble component; at the upper level, the ensemble maintains a shared beam by voting and synchronizing the candidates (sub-sequences) in individual beams.
-Unlike word-level ensembles Freitag et al.
-(2017); Shanbhogue et al.
-(2023), EBBS does not average the predicted distributions, encouraging individual predictors to explore their own preferences; unlike sequence-level ensembles Kobayashi (2018); Eikema and Aziz (2020), EBBS does not select from a candidate set, and thus is more flexible since voting is performed throughout the generation process.
-We conducted experiments on IWSLT (Cettolo et al., 2017) and Europarl (Koehn, 2005), two popular multilingual datasets for zero-shot machine translation.
+Our EBBS performs two levels of beam search at each generation step: at the lower level, beam search is applied individually to each ensemble component; at the upper level, the ensemble maintains a shared beam by voting and synchronizing the candidates (sub-sequences) in lower-level beams.
+Unlike word-level ensembles (Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al.
+2023), EBBS does not average the predicted distributions, encouraging individual predictors to explore their own preferences; unlike sequence-level MBR ensembles (Kobayashi 2018; Eikema and Aziz 2020), EBBS does not select from a candidate set, and thus is more flexible since votings are performed throughout the generation process.
+We conducted experiments on IWSLT (Cettolo et al.
+2017) and Europarl (Koehn 2005), two popular multilingual datasets for zero-shot machine translation.
 Results show that EBBS can generate high-quality translations and outperform existing ensemble techniques.
-In addition, we used EBBS-generated outputs as the distillation signal to further finetune our multilingual translation model, which largely improves the efficiency while retaining the performance or even surpassing EBBS on Europarl.
+In addition, we used EBBS-generated data for distillation to further improve the multilingual model.
+The experiment shows that such a distilling process encourages the model to learn from high-quality translations produced by EBBS, allowing it to outperform EBBS with no inference overhead compared with direct translation.
 2 Related Work
-Zero-shot machine translation.
-With monolingual data, it is possible to accomplish zero-shot machine translation by training language-specific autoencoders that share the same latent space (Lample et al., 2018a, b; Mohiuddin and Joty, 2020).
-In a multilingual setting, zero-shot ability emerges when a single model is trained with multiple translation directions, allowing it to generalize to unseen directions (Dabre et al., 2020; Ranathunga et al., 2023; Johnson et al., 2017; Fan et al., 2021).
-A key research direction for multilingual translation is to alleviate task interference among different translation directions (Zaremoodi et al., 2018; Wang et al., 2020).
-Researchers have developed various parameter-sharing schemes Baziotis et al.
-(2022); Chronopoulou et al.
-(2023) and applied optimization-based methods to locate language-specific neurons Wang and Zhang (2022); He et al.
-(2023); such a model is able to better learn individual translation directions while pooling the knowledge of all languages together.
+Machine translation.
+In NLP, machine translation is a long-standing task that aims to rewrite text from one language to another without changing the meaning.
+Traditional research in translation has been mainly centered on the supervised setting, utilizing manually crafted rules (Forcada et al.
+2011; Dugast, Senellart, and Koehn 2007) and statistical methods (Brown et al.
+1990; Koehn 2009); more recently, neural machine translation systems have considerably improved the performance (Vaswani et al.
+2017; Raffel et al.
+2020).
+However, translation remains challenging for low-resource languages, where neural models do not have enough parallel data to train on.
+Translation for low-resource languages largely relies on zero-shot techniques, where no parallel text is available for a particular translation direction.
+In general, zero-shot translation can be accomplished in a monolingual or multilingual setting.
+With monolingual data, the most common approach is to build language-specific autoencoders that share the same latent space of semantics; translation is then achieved by plugging in the decoder of the desired language (Lample et al.
+2018a, b; Mohiuddin and Joty 2020).
+In this paper, we focus on the multilingual setting, where one model can translate between multiple languages (Dabre, Chu, and Kunchukuttan 2020).
+Usually, parallel texts only exist for a high-resource language such as English, leaving translations between low-resource languages zero-shot (e.g., Italian to Dutch) (Johnson et al.
+2017; Fan et al.
+2021).
+In this setting, the most common approach is to train the multilingual model on English-centric data, and the zero-shot translation ability naturally emerges during the training process (Johnson et al.
+2017; Scao et al.
+2022).
+A key challenge for multilingual models is task interference, where too many languages tend to degrade model performance (Zaremoodi, Buntine, and Haffari 2018; Wang, Lipton, and Tsvetkov 2020).
+As a result, research in this direction has been alleviating such interference by developing various parameter-separation schemes (Baziotis et al.
+2022; Chronopoulou, Stojanovski, and Fraser 2023) and using gradient-based methods to update language-specific parameters (Wang and Zhang 2022; He et al.
+2023).
 In our work, we use a standard Transformer model following Johnson et al.
@@
 (2021).
-Our proposed ensemble algorithm EBBS is compatible with the above methods, as it is agnostic to model architectures.
-Ensemble learning.
-Ensemble methods integrate multiple machine learning systems to form a stronger one (Dong et al., 2020; Yang et al., 2023).
-Bagging, a classic ensemble technique, works by training multiple models with different portions of data and combining their predictions by averaging or voting (Breiman, 1996; Bühlmann and Yu, 2002).
-Another popular ensemble approach is boosting, where multiple models are trained sequentially, with each subsequent model focusing on the mistakes of the previous ones (Schapire, 2003; Hastie et al., 2009; Natekin and Knoll, 2013).
-Unfortunately, bagging and boosting are not compatible with our setting, because we build a self-ensemble with only a single model.
-Alternatively, stacking combines the outputs by training a meta-model (Wolpert, 1992; Ganaie et al., 2022), but this does not apply in our zero-shot setting either because we do not have groundtruth signal to train the meta-model.
-These ensemble techniques nevertheless may be applied to supervised text generation (Freitag et al., 2017; Kobayashi, 2018; Hendy et al., 2021).
-That being said, they are not ideal as they do not take advantage of the sequential nature of sentences.
-Unlike previous work, our EBBS performs bi-level beam search, allowing each component to explore its own predictions at every step but synchronizing all individuals by a “soft voting” mechanism.
+Our proposed ensemble algorithm EBBS is compatible with the above approaches, as it is agnostic to model architectures.
+Ensemble methods.
+In a model ensemble, multiple machine learning systems are integrated so as to form a stronger one (Dong et al.
+2020; Yang, Lv, and Chen 2023).
+Bagging, a classic ensemble technique, works by training multiple models with different portions of data and combining their predictions through averaging or voting (Breiman 1996; Bühlmann and Yu 2002).
+Another popular ensemble approach is boosting, where different models are trained sequentially, with each subsequent model addressing the mistakes of the previous ones (Schapire 2003; Hastie et al.
+2009; Natekin and Knoll 2013).
+Unfortunately, bagging and boosting are not compatible with our setting, because we build an ensemble with a single model.
+Alternatively, stacking combines the outputs by training a meta-model (Wolpert 1992; Ganaie et al.
+2022), but this does not apply to our zero-shot setting either because we do not have groundtruth signals to train the meta-model.
+Even though these ensemble techniques may be directly applied to supervised generation (Freitag, Al-Onaizan, and Sankaran 2017; Kobayashi 2018; Hendy et al.
+2021), they are not ideal as they do not take advantage of structured prediction.
+Our recent work has addressed the ensemble of tree structures (Shayegh et al.
+2024; Shayegh, Wen, and Mou 2024; Shayegh et al.
+2025), and in this paper we focus on text generation.
+Unlike previous work, our EBBS performs bi-level beam search, exploring different components’ own predictions and synchronizing them by a “soft voting” mechanism at every step.
 Our approach is specifically suited to the sequence generation process.
@@
 In this section, we first explain our ensemble components in §3.1.
-We then propose EBBS, a novel ensemble decoding algorithm in §3.2.
-Finally, we describe in §3.3 knowledge distillation with EBBS-decoded outputs to improve efficiency.
+In §3.2, we propose EBBS, a novel ensemble decoding algorithm.
+Finally, we describe in §3.3 knowledge distillation with EBBS-decoded outputs for efficiency considerations.
 3.1 Ensemble Components
 In this work, we focus on zero-shot multilingual machine translation, which requires a system to perform translations for multiple languages, where some translation directions are unseen.
-Specifically, our multilingual model is an encoder–decoder Transformer with a byte pair encoding tokenizer (Sennrich et al., 2016b) shared among all languages.
-The encoder can capture the semantics of tokens in any language, whereas the decoder translates the encoded text into the desired language based on a target-language indicator Johnson et al.
-(2017); Fan et al.
-(2021).
-We follow the standard English-centric training (Johnson et al., 2017; Liu et al., 2021), where the multilingual model is trained with parallel data with English on one of the sides (e.g., German-to-English and English-to-Romanian).
-The zero-shot ability emerges during such training, and the model is able to translate directly between unseen language pairs (e.g., German-to-Romanian), known as direct translation Dabre et al.
-(2020); Ranathunga et al.
-(2023).
+Specifically, our multilingual model is an encoder–decoder Transformer with a byte pair encoding tokenizer (Sennrich, Haddow, and Birch 2016b) shared among all languages.
+The encoder can capture the semantics of tokens in different languages, whereas the decoder translates the encoded text into the desired language based on a target-language indicator token (Johnson et al.
+2017; Fan et al.
+2021).
+We follow the standard English-centric training (Johnson et al.
+2017; Liu et al.
+2021), where the multilingual model is trained using parallel data with English on one side (e.g., German-to-English and English-to-Romanian).
+As mentioned in §1, the zero-shot ability emerges during such training, and the model is able to perform direct translation between unseen language pairs (e.g., German-to-Romanian) (Dabre, Chu, and Kunchukuttan 2020; Ranathunga et al.
+2023).
 An alternative approach is pivot translation, where the multilingual model performs two translations using a high-resource language as a pivot (e.g., first translating German to English, and then English to Romanian).
-However, both direct and pivot translations have major weaknesses: the quality of direct translation tends to be low due to the lack of parallel data, whereas pivot translation suffers from an error accumulation as it requires two translation steps (Babych et al., 2007; Gu et al., 2019).
-In this paper, we would like to build a self-ensemble of the multilingual model to boost translation quality, where direct and pivot translations are our ensemble components.
-However, standard ensemble methods such as averaging and voting may not work well for text generation.
+However, both direct and pivot translations have major weaknesses: the quality of direct translation tends to be low due to the lack of parallel data, whereas pivot translation suffers from error accumulation as it requires two translation steps (Babych, Hartley, and Sharoff 2007; Gu et al.
+2019).
+In this paper, we would like to build an ensemble of direct and pivot translations to boost translation quality, where each translation path results in an ensemble component.
+Commonly used ensemble methods such as averaging and voting may not work well for text generation.
 Voting, for example, chooses the most voted prediction, but in text generation, the components’ votes often do not share anything in common, because there could be tens of thousands of tokens in the vocabulary.
 An averaging ensemble, on the other hand, averages the predicted distributions of all components, potentially leading to an overly smooth distribution.
-As reported in a pilot study in the appendix of Fan et al.
-(2021), a multilingual averaging ensemble only achieves a small improvement over direct translation.
-Overall, ensemble methods have not been widely explored for zero-shot multilingual translation in previous literature.
+Despite early success by Razmara and Sarkar (2013), more recent studies report marginal or negative improvement for multi-pivot averaging ensemble (Fan et al.
+2021; Gaikwad et al.
+2024; Mohammadshahi, Vamvas, and Sennrich 2024).
 3.2 Our Proposed EBBS Algorithm
-Figure 1: Overview of our EBBS algorithm.
+Figure 1: Illustration of our EBBS algorithm.
+We propose an ensemble with bi-level beam search (EBBS), a novel decoding algorithm that enables different ensemble components to collaborate and vote on each other’s partial generations with two levels of beam search.
+At the lower level, each ensemble component performs beam search individually, exploring its own preferred regions of the sentence space.
+At the upper level, EBBS synchronizes the lower-level beam candidates through a voting mechanism, only keeping the most promising partial generations in a shared, upper-level beam.
+This allows the ensemble components to vote out spurious partial candidates and improve zero-shot translation performance.
+Concretely, we assume there are $K$ ensemble components $p_{1},\cdots,p_{K}$, each predicting the probability of the next word given some prefix.
+For the 00th decoding step, EBBS initializes the upper-level beam by $\overline{B}_{0}=\langle\textsc{bos},1\rangle$, suggesting that a sequence is forced to start with a special beginning-of-sequence token bos with probability $1$.
+For step $t$, each ensemble component performs lower-level beam search individually, based on the prefixes in the last step’s shared beam $\overline{B}_{t-1}$:
+[…]
+$\langle\mathbf{y}_{1:t-1},p\rangle\in\overline{B}_{t-1},\ \ \mathrm{y}\in V\ \}$ (1)
+for $k=1,\cdots,K$.
+Here, $\operatorname{top-}\!Z$ selects $Z$-many sequences with the highest probabilities, $\oplus$ represents string concatenation, $V$ is the vocabulary, and $p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$ is the $k$th ensemble component’s predicted probability at step $t$ given the prefix $\mathbf{y}_{1:t-1}$ and input $\mathbf{x}$.
+At the upper level, EBBS synchronizes the lower-level individual beams $\underline{B}_{t,k}$, for $k=1,\cdots,K$, into a shared, upper-level beam through a soft-voting mechanism, where the candidate set $C_{t}$ is the union of the sequences in lower-level beams:
+$C_{t}=\bigcup\nolimits_{k}\{\mathbf{y}:\langle\mathbf{y},p\rangle\in\underline{B}_{t,k}\}$ (2)
+We evaluate each candidate in $C_{t}$ and compute its overall vote as the sum of the probabilities.
+[…]
+In this way, the upper level synchronizes all ensemble components with the shared beam $\overline{B}_{t}$ for the next step of generation.
+Intuitively, our voting scheme gives an ensemble component $Z$-many votes, each weighted by the predicted probability.
+The votes (probabilities) are then tallied (summed) for each candidate to form the upper-level beam.
+Our bi-level beam search terminates when we have $Z$-many terminated sequences in the shared beam, and returns the sequence with the highest score [^1] For selecting the final output, we follow standard implementations and normalize the joint probability by length, i.e., taking the geometric mean of step-wise probabilities (Wolf et al.
+2019; Ott et al.
+2019).
+Otherwise, beam search algorithms are often biased towards short sequences (Meister, Cotterell, and Vieira 2020).
+as the ensemble output.
+We provide the detailed pseudocode for EBBS in Algorithm 1 and an illustration in Figure 1.
+Discussion.
+Traditional beam search keeps a fixed-size beam of high-likelihood partial sequences.
+To build an ensemble with multiple predictors, it is tempting to directly average their probabilities $p(\mathbf{y}|\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}p_{k}(\mathbf{y}|\mathbf{x})$ as the score for beam search, which has been experimented in previous work (Sennrich, Haddow, and Birch 2016a; Shanbhogue et al.
+2023).
+However, our intuition suggests that such an approach may suffer from the over-smoothing problem (Wei et al.
+2019; Wen et al.
+2023b): when multiple translations (known as modes) are plausible given an input, the ensemble process will overly smooth out the modes by probability averaging.
+By contrast, EBBS allows each ensemble component to explore its own mode (Lines 4–11, Algorithm 1).
+In Figure 1, for example, the top sequence yields two plausible next tokens, suggested by each component in the lower level; their probabilities are not smoothed out in our approach, unlike averaging ensembles.
+The upper level performs soft voting (Lines 12–19, Algorithm 1) so as to maintain tractable inference.
+3.3 EBBS-Based Distillation
 Input: $\mathbf{x}$: input sentence; $Z$: beam size
@@
 1 $H\leftarrow\emptyset$triangleright candidate outputs
-2 $B_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright upper-level beam
+2 $\overline{B}_{0}\leftarrow\{\langle\textsc{bos},1\rangle\}$triangleright upper-level beam
 3 for $t=1,2,\cdots$ do
-4       $\triangleright$ lower level: individual beam search
-5       for $\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1}$ do
+4       $\triangleright$ lower: individual beam search
+5       for $\langle\mathbf{y}_{1:t-1},p\rangle\in\overline{B}_{t-1}$ do
 6             for $k=1,\cdots,K$ do
-7                   $B^{\prime}_{t,k}\leftarrow\emptyset$triangleright lower-level beam
+7                   $\underline{B}_{t,k}\leftarrow\emptyset$triangleright lower-level beam
 8                   for $\mathrm{y}\in V$ do
 9                         $p^{\prime}\leftarrow p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$
-10                         $B^{\prime}_{k}.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p\cdot p^{\prime}\rangle)$
-11                  $B^{\prime}_{t,k}\leftarrow B^{\prime}_{t,k}.\operatorname{top}(Z)$
+10                         $\underline{B}_{t,k}.\operatorname{add}(\langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},p\cdot p^{\prime}\rangle)$
+11                  $\underline{B}_{t,k}\leftarrow\underline{B}_{t,k}.\operatorname{top}(Z)$
 12
-13      $\triangleright$ upper level: beam synchronization
+13      $\triangleright$ upper: beam synchronization
 14       $D\leftarrow\text{empty dictionary}$
 15       for $k=1,\cdots,K$ do
-16             for $\langle\mathbf{y},p\rangle\in B^{\prime}_{t,k}.\operatorname{top}(Z)$ do
+16             for $\langle\mathbf{y},p\rangle\in\underline{B}_{t,k}$ do
 17                   if $\mathbf{y}\in D$ then
@@
 22
-23      $B_{t}\leftarrow D.\operatorname{top}(Z)$
+23      $\overline{B}_{t}\leftarrow D.\operatorname{top}(Z)$
 24       $\triangleright$ check for termination
-25       for $\langle\mathbf{y},p\rangle\in B_{t}$ do
+25       for $\langle\mathbf{y},p\rangle\in\underline{B}_{t}$ do
 26             if $\mathrm{y}_{t}=\textsc{eos}$ then
@@
 Algorithm 1 Our EBBS Algorithm
-To this end, we propose an ensemble with bi-level beam search (EBBS), a novel decoding framework that enables different ensemble components to collaborate and vote on each other’s partial generations with two levels of beam search.
-At the lower level, each ensemble component performs beam search individually, exploring its own preferred regions of the sentence space.
-At the upper level, EBBS synchronizes the lower-level beam candidates through a voting mechanism, only keeping the most promising partial generations in a shared, upper-level beam.
-This allows the ensemble components to vote out spurious partial candidates and improve zero-shot translation performance.
-Concretely, we assume there are $K$ ensemble components $p_{1},\cdots,p_{K}$, each predicting the probability of the next word given some prefix.
-For the 00th decoding step, EBBS initializes a shared, upper-level beam $B_{0}=\langle\textsc{bos},1\rangle$, suggesting that a sequence is forced to start with a special token bos with probability $1$.
-For step $t$, each ensemble component performs a lower-level beam search individually, based on the prefixes in the last step’s shared beam $B_{t-1}$:
-$B_{t,k}^{\prime}=$ $\operatorname{top-}\!Z\{\ \langle\mathbf{y}_{1:t-1}\oplus\mathrm{y},\ \ p\cdot p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})\rangle:$
-$\hskip 37.5807pt\langle\mathbf{y}_{1:t-1},p\rangle\in B_{t-1},\ \ \mathrm{y}\in V\}$ (1)
-for $k=1,\cdots,K$.
-Here, $\operatorname{top-}\!Z$ selects $Z$-many sequences with the highest probabilities, $\oplus$ represents string concatenation, $V$ is the vocabulary, and $p_{k}(\mathrm{y}|\mathbf{y}_{1:t-1},\mathbf{x})$ is the $k$th ensemble component’s predicted probability at step $t$ given the prefix $\mathbf{y}_{1:t-1}$ and input $\mathbf{x}$.
-At the upper level, EBBS synchronizes the individual beams $B_{t,k}^{\prime}$, for $k=1,\cdots,K$, into a single, shared, upper-level beam through a soft voting mechanism, where the candidate set $C_{t}$ is the union of the sequences in individual beams:
-$C_{t}=\bigcup\nolimits_{k}\{\mathbf{y}:\langle\mathbf{y},p\rangle\in B_{t,k}^{\prime}\}$ (2)
-Then, we evaluate each candidate in $C_{t}$ and compute its overall vote as the sum of the probabilities.
-[…]
-In this way, the upper level synchronizes every ensemble component with the shared beam $B_{t}$ for the next step of generation.
-Intuitively, our voting scheme gives an ensemble component $Z$-many votes, each weighted by the predicted probability.
-The votes (probabilities) are then tallied (summed) for each candidate to form the upper-level beam.
-Our bi-level beam search terminates when we have $Z$-many terminated sequences in the shared beam, and returns the sequence with the highest score [^1] For selecting the final output, we follow standard implementations and normalize the joint probabilities by length, i.e., taking the geometric mean of step-wise probabilities (Wolf et al., 2019; Ott et al., 2019).
-Otherwise, beam search algorithms are often biased towards short sequences.
-as the ensemble output.
-We provide the detailed pseudocode for EBBS in Algorithm 1 and an illustration in Figure 1.
-Discussion.
-As shown in Algorithm 2 (Appendix A), traditional beam search keeps a fixed-size beam of high-likelihood partial sequences.
-It is tempting to combine such vanilla beam search with multiple predictors by directly averaging their probabilities $p(\mathbf{y}|\mathbf{x})=\frac{1}{K}\sum_{k=1}^{K}p_{k}(\mathbf{y}|\mathbf{x})$ as the score for beam search (Line 8, Algorithm 2), which has been experimented in previous work Sennrich et al.
-(2016a); Shanbhogue et al.
-(2023).
-However, our intuition suggests that such an approach may suffer from the over-smoothing problem Wei et al.
-(2019); Wen et al.
-(2023b): when multiple translations (known as modes) are plausible given an input, the ensemble process will overly smooth out the modes by probability averaging.
-By contrast, EBBS allows each ensemble component to explore its own mode (Lines 4–11, Algorithm 1).
-In Figure 1, for example, the top sequence yields two plausible next tokens, suggested by each component in the lower level.
-Their probabilities are not smoothed out in our approach, unlike averaging ensembles.
-The upper level performs soft voting (Lines 12–19, Algorithm 1) so as to maintain tractable inference.
-3.3 EBBS-Based Distillation
-To improve inference efficiency, we may perform knowledge distillation based on the outputs of our EBBS algorithm.
-In particular, we follow Kim and Rush (2016) and apply a sequence-level knowledge distillation loss, treating the output $\hat{\mathbf{y}}$ of our ensemble (called a teacher) as the pseudo-groundtruth for finetuning the multilingual translation model (called a student):
+To improve inference efficiency, we perform knowledge distillation based on the outputs of our EBBS algorithm.
+In particular, we follow (Kim and Rush 2016) and apply a sequence-level knowledge distillation loss, treating the output $\hat{\mathbf{y}}$ of our ensemble (serving as a teacher) as the pseudo-groundtruth for finetuning the multilingual translation model (serving as a student):
 $\mathcal{L}_{\text{KD}}=-\sum_{t=1}^{|\hat{\mathbf{y}}|}\log p(\hat{\mathrm{y}}_{t}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})$ (4)
-Our self-distillation is an ensemble-then-distill process.
-This differs from a straightforward practice of multi-teacher distillation, where the student learns from the union of teachers’ outputs Wu et al.
-(2021).
-The commonly applied cross-entropy loss is known to yield overly smooth distributions Wen et al.
-(2023a, b), and the problem becomes more severe with multiple teachers, leading to unsatisfactory performance of union distillation Shayegh et al.
-(2024).
-By contrast, our approach provides the student with a consolidated pseudo-groundtruth translation for an input, causing less confusion during the distillation process especially when teachers disagree with each other.
+Our distilling method is an ensemble-then-distill process.
+This differs from a straightforward practice of multi-teacher distillation, where the student learns from the union of teachers’ outputs (Wu, Wu, and Huang 2021).
+The commonly applied cross-entropy loss is known to yield overly smooth distributions (Wen et al.
+2023a, b), and the problem becomes more severe with multiple teachers, leading to less satisfactory performance of union distillation (Shayegh et al.
+2024).
+On the contrary, our approach provides the student with a consolidated pseudo-groundtruth translation, causing less confusion during the distillation process especially when teachers disagree.
 4 Experiments
 4.1 Settings
-Datasets.
-We evaluated EBBS on two popular benchmark datasets for zero-shot machine translation: IWSLT (Cettolo et al., 2017), which contains 4 languages (with English) and 6 zero-shot directions; and Europarl v7 (Koehn, 2005), which contains 9 languages and 56 zero-shot directions.
+We evaluated EBBS on two popular benchmark datasets for zero-shot machine translation: IWSLT (Cettolo et al.
+2017), which contains 4 languages (with English) and 6 zero-shot directions; and Europarl v7 (Koehn 2005), which contains 9 languages and 56 zero-shot directions.
 […]
-Model architecture.
-We replicated Liu et al.
-(2021) and trained a multilingual system as our base model.
-Specifically, we had a 5-layer encoder–decoder Transformer for IWSLT, but used an 8-layer Transformer for Europarl to accommodate more training data and languages.
-We also followed Liu et al.
-(2021) and removed residual connections in certain middle layers of the encoder (2nd layer for IWSLT; 2nd and 4th layers for Europarl).
-Liu et al.
-(2021) show weakening positional correspondence between layers can significantly improve zero-shot performance, which was also observed in our replication experiments.
-More details about our experimental setup can be found in Appendix B.
+We replicated (Liu et al.
+2021) and trained a multilingual translation system as our base model.
+Specifically, the neural architecture in (Liu et al.
+2021) is a 5-layer encoder–decoder Transformer for IWSLT, but has 8 layers for Europarl to accommodate more training data and languages.
+For EBBS, we used a beam size of five for both upper- and lower-level beams.
+In our experiment, we implemented standard beam search for comparison, where we also used a beam size of five, following the common practice (Meister, Cotterell, and Vieira 2020).
+A comprehensive beam analysis can be found in our appendix.
 4.2 Competing Methods
-We compare our EBBS against the following methods, including pivoting and different ensembles.
-English pivoting.
-Pivoting is commonly applied for translation (Wu and Wang, 2007, 2009; Vamvas and Sennrich, 2022), and we used English as the pivot in our setting because we have parallel data for translations both from and to English.
+We comprehensively compare our EBBS with direct/pivot translation and other ensemble methods.
+Direct/pivot translation.
+For direct translation, we applied beam search on the multilingual model to translate in unseen directions.
+For pivot translation (Wu and Wang 2007, 2009; Vamvas and Sennrich 2022), we used English as the pivot because we have parallel data for translations both from and to English.
 Word-level averaging ensemble.
-Averaging is one of the most widely used ensemble techniques in text generation (Sennrich et al., 2016a; Freitag et al., 2017; Shanbhogue et al., 2023).
-Essentially, the ensemble components’ probabilities are first averaged before being fed to the standard beam search (Algorithm 2).
+Averaging is one of the most widely used ensemble techniques in text generation (Sennrich, Haddow, and Birch 2016a; Freitag, Al-Onaizan, and Sankaran 2017; Shanbhogue et al.
+2023).
+Essentially, the ensemble components’ probabilities are first averaged before being fed to the standard beam search.
 Word-level voting ensemble.
@@
 Sequence-level voting ensemble.
-Minimum Bayes risk decoding is originally designed as a single-model decoding algorithm, where it selects a sequence from a set of beam search results based on similarity (Eikema and Aziz, 2020; Müller and Sennrich, 2021).
+Minimum Bayes risk (MBR) decoding is originally designed as a single-model decoding algorithm, where it selects a sequence from a set of beam search results based on similarity (Eikema and Aziz 2020; Müller and Sennrich 2021).
 Here, we use it as a sequence-level ensemble technique, where the candidates are the output sequences from different ensemble components.
@@
 4.3 Results and Analysis
-IWSLT # Model Average it-nl it-ro nl-it nl-ro ro-it ro-nl
-1 Direct translation (Liu et al., ${}^{\dagger}$ 17.7 18.5 17.8 17.9 15.5 19.6 16.8
-[…]
+IWSLT # Method Average it-nl it-ro nl-it nl-ro ro-it ro-nl
+1 Direct translation (Liu et al. 2021) 17.7 18.5 17.8 17.9 15.5 19.6 16.8
+2 Direct translation (our replication) 17.29 17.46 17.48 18.23 14.63 19.65 16.26
 3 Pivoting (en) 16.19 17.49 15.09 16.79 13.05 18.34 16.37
-4 Word-level averaging ensemble 16.52 16.48 16.49 17.53 13.80 19.07 15.77
-[…]
-6 Sequence-level voting ensemble (MBR) 16.66 16.54 16.51 17.72 13.64 19.58 15.98
-[…]
-Europarl # Model Average da-de da-es da-fi da-fr da-it da-nl da-pt
-1 Direct translation (Liu et al., ${}^{\dagger}$ 26.9 24.2 33.1 18.1 30.6 26.1 26.3 29.9
-[…]
-[…]
-4 Word-level averaging ensemble 27.76 26.21 33.64 18.88 31.09 26.71 27.40 30.37
-[…]
-[…]
+4 Word-level averaging ensemble 17.28 17.29 17.44 18.33 14.65 19.69 16.30
+5 Word-level voting ensemble 16.99 17.58 16.38 17.78 14.13 19.21 16.84
+6 Sequence-level voting ensemble (MBR) 16.72 16.64 16.53 17.83 13.74 19.48 16.08
+7 EBBS (ours) 18.24 19.52 17.09 19.06 14.58 20.75 18.45
+8 Direct w/ EBBS distillation (ours) 18.92 19.86 18.80 19.73 15.39 21.23 18.48
+Europarl # Method Average da-de da-es da-fi da-fr da-it da-nl da-pt
+1 Direct translation (Liu et al. 2021) 26.9 24.2 33.1 18.1 30.6 26.1 26.3 29.9
+2 Direct translation (our replication) 27.74 26.24 33.64 18.95 31.01 26.58 27.36 30.38
+3 Pivoting (en) 27.69 25.17 33.87 18.70 31.44 27.12 26.75 30.79
+4 Word-level averaging ensemble 27.76 26.13 33.72 18.91 31.01 26.67 27.39 30.50
+5 Word-level voting ensemble 27.45 25.76 33.24 18.39 30.96 26.83 26.63 30.37
+6 Sequence-level voting ensemble (MBR) 27.90 25.90 33.95 19.15 31.50 27.15 27.09 30.55
 7 EBBS (ours) 28.36 26.32 34.28 19.43 31.97 27.67 27.78 31.08
-Table 1: Main results on the IWSLT and Europarl datasets.
-The best results are in bold, and the second best results are underlined.
-${}^{\dagger}$ indicates cited results; others were obtained by our experimentation.
+8 Direct w/ EBBS distillation (ours) 28.54 26.75 34.68 19.89 32.00 27.69 27.61 31.19
+Table 1: Main results of BLEU scores on IWSLT and Europarl. The best results are in bold; the second best results are underlined. indicates cited results; others were obtained by our experimentation.
 Main results.
-We first replicated the base multilingual model Liu et al.
-(2021).
-As shown in Rows 1–2, Table 1, the results are generally close, indicating that the replication is successful and ready for our ensemble research.
+Our experiment starts by a replication of the base multilingual model (Liu et al.
+2021).
+As shown in Rows 1–2, Table 1, the results are generally close, indicating that our replication is successful and ready for ensemble research.
 Further, we tried English pivoting (Row 3), a common zero-shot translation method.
-In our experiments, we find that it does not outperform direct translation, as pivoting methods may suffer from the error accumulation problem since it requires two steps of translation.
-We then compare different ensemble techniques, including the proposed EBBS and alternative ensemble methods.
-Specifically, we notice that IWSLT contains four languages (with English); thus we have two available pivoting directions (excluding source and target), which, along with direct translation, are our three ensemble components.
+In our experiments, we find that it does not outperform direct translation, as pivoting methods may suffer from the error accumulation problem due to two-step translation.
+We then compare different ensemble techniques, including our proposed EBBS.
+We notice that IWSLT contains four languages (with English); thus we have two available pivoting directions (excluding source and target), which, along with direct translation, are our three ensemble components.
 For Europarl, it contains nine languages; for performance and efficiency concerns (to be shown in Figure 2), we also consider three translation paths as our ensemble components: direction translation, English pivoting, and a second pivot.
@@
 For example, Spanish-to-German translation will have to use French as the pivot.
-These languages are chosen because they have the most content on the Internet according to the Web Technology Surveys (https://w3techs.com/technologies/overview/content_language), allowing us to better simulate real-world applications.
-We start by studying the standard ensemble technique of word-level averaging (Row 4), which has been used in previous translation research (Freitag et al., 2017).
-In particular, our averaging ensemble performs worse than direct translation on IWSLT, but is slightly better on Europarl.
-Our results are not consistent to Freitag et al.
-(2017), who show a word-level averaging ensemble of random seeds can improve performance.
-This is understandable because models trained with random seeds exhibit similar behavior, and averaging their probabilities achieves a denoising effect.
-However, our ensemble components are drastically different in terms of their strengths and expertise, because they learn diverse translation patterns from different languages.
+These languages are chosen because they have the most content on the Internet according to the Web Technology Surveys (https://w3techs.com/technologies/overview/content˙language).
+We study the common ensemble technique of word-level averaging (Row 4), which has been used in previous translation research (Freitag, Al-Onaizan, and Sankaran 2017).
+As we can see, the averaging ensemble performs similarly to direct translation on both datasets.
+Our zero-shot results are different from (Freitag, Al-Onaizan, and Sankaran 2017), which shows a word-level averaging ensemble of random seeds can improve performance in the supervised setting.
+This is because models trained with different random seeds exhibit similar behavior, and averaging their probabilities achieves a denoising effect.
+However, our ensemble components differ drastically in terms of their strengths and expertise due to the different translation paths (direct and pivot translations).
 Thus, word averaging fails to improve translation quality in our setting.
 Alternatively, voting ensembles can also be applied, at either the word level or the sequence level.
-As seen, word-level voting is not effective, as it does not achieve significant improvements (Row 5).
+As seen, word-level voting is not effective, as it is worse than direct translation on both datasets (Row 5).
 This is expected because the voted words (top predictions) by the ensemble components may not overlap due to the large vocabulary size.
@@
 Sequence-level voting should also be done in a soft manner, and minimum Bayes risk (MBR) decoding can be thought of as using a Bayes risk to softly “vote” the candidate outputs.
-As seen from Row 6, such a method works relatively well on Europarl, achieving the second-highest performance across all competing methods; however, it works poorly on the IWSLT dataset.
-The main drawback of sequence-level voting is that it can only select one of the ensemble components’ outputs.
-This may not work well when the individual ensemble components are weak, as in our zero-shot setting, and especially with the IWSLT dataset due to a relatively small training set.
+As seen from Row 6, such a method works relatively well on Europarl, achieving the second-highest performance across all ensemble methods; however, it works poorly on the IWSLT dataset.
+The main drawback of sequence-level voting is that it can only select one of the ensemble components’ output.
+This may not work well when the individual ensemble components are weak, especially with the small IWSLT dataset.
 Such a selective sequence-level ensemble cannot integrate different expertise of its components during generation.
 Dataset Method Avg. BLEU Wins Losses
-IWSLT Direct 17.28 2 4
-Ensemble 18.23 4 2
-Europarl Direct 27.85 4 52
-Ensemble 28.44 52 4
-Overall Direct 26.83 6 56
-Ensemble 27.45 56 6
+IWSLT Direct translation 17.29 2 4
+EBBS (ours) 18.24 4 2
+Europarl Direct translation 27.85 4 52
+EBBS (ours) 28.44 52 4
+Overall Direct translation 26.83 6 56
+EBBS (ours) 27.45 56 6
 $p$-value 3e-11
-Table 2: Performance statistics on all 66 zero-shot directions in both datasets. $p$-value is given by a two-sided binomial test.
-Unlike existing ensemble methods, our EBBS algorithm achieves the highest performance in most directions on both datasets.
-Notice that Europarl contains 56 zero-shot directions; due to the space limit, we can only present in Table 1 the first seven directions based on the order provided by the dataset.
+Table 2: Pairwise comparison on all 62 zero-shot directions in both datasets. The $p$-value is given by a two-sided binomial test.
+Unlike existing ensemble methods, our EBBS algorithm achieves higher performance in most directions on both datasets.
+Noticing that Europarl contains 56 zero-shot directions, we could only present in Table 1 the first seven directions based on the order provided by the dataset, due to the space limit.
 Table 2 further shows a pairwise comparison against direct translation (a strong baseline in our experiment) in all zero-shot directions.
 As seen, EBBS achieves higher performance in 56 out of 62 cases across two datasets, showing strong statistical evidence for its effectiveness, with a $p$-value of 3e-11 in a two-sided binomial test.
-Voting scheme BLEU BLEU1 BLEU2 BLEU3 BLEU4
-None (beam search) 25.33 56.32 30.08 19.01 12.78
-Total-sum 25.27 56.67 30.27 19.07 12.65
-[…]
-[…]
-Top-$Z$ sum (ours) 26.10 57.07 31.00 19.76 13.28
-Table 3: Comparison of different ensemble variants. We used Italian-to-Dutch translation in the Europarl dataset as a testbed due to the limit of time and resources.
-Analysis of voting methods in EBBS.
-In our EBBS algorithm, the lower-level beams are synchronized into a shared upper-level beam by voting.
-Specifically, EBBS uses a sum-voting mechanism, where we add the ensemble components’ probabilities for each appearance of a candidate in the lower-level beam, shown in Eqn.
-(3).
-Here, we analyze a few alternative voting methods for EBBS.
-If EBBS adopts total-sum voting, it still uses lower-level beams to find candidates, but adds all components’ probabilities together.
-This is equivalent to applying the common averaging ensemble to the top-$Z$ candidates.
-However, it differs from our approach, because in total-sum voting, a component will vote even if the candidate does not appear in its own lower-level beam; the probability after voting in Eqn.
-(3) is substituted with $\frac{1}{K}\sum_{k}p_{k}(\mathbf{y}|\mathbf{x})$.
-As shown in Table 3, EBBS with total-sum voting performs worse than direct translation, suggesting the importance of ignoring the components whose lower-level beam does not contain the candidate.
-This is analogous to nucleus sampling Holtzman et al.
-(2019), where the long tail of a distribution is mainly noise and should be ignored.
-Other voting schemes that EBBS may use include 0/1 voting and max voting.
-The former selects the candidates that appear most in the lower-level beams, disregarding the probability values (unless for ties); the latter chooses the maximum probability across the lower-level beams, which gives preference to sequences through a maximization bias (Hasselt, 2010; van Hasselt et al., 2016).
-As seen in Table 3, EBBS performs relatively well with both of these voting schemes, achieving a decent improvement over the baseline approach; however, their performance is consistently worse than our sum voting scheme.
-Overall, our bi-level beam search ensemble is effective with different voting schemes (except for the total-sum voting), and our sum voting works the best among these variants.
-Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
-Direct 25.33 56.32 30.08 19.01 12.78
-Pivoting (en) 25.08 56.76 30.29 19.06 12.66
-Pivoting (es) 24.40 55.38 29.08 18.22 12.09
-Pivoting (pt) 24.34 55.46 29.02 18.13 12.02
-Pivoting (fr) 24.20 55.41 29.02 18.00 11.84
-Pivoting (de) 23.65 55.33 28.69 17.67 11.54
-Pivoting (da) 23.12 54.81 27.96 17.12 11.12
-Pivoting (fi) 20.74 53.54 26.10 15.43 9.79
-[…]
-Top-3 ensemble 26.10 57.07 31.00 19.76 13.28
-Table 4: The performance of individuals and their ensembles for Italian-to-Dutch translation on Europarl.
-[…]
-Notably, EBBS can achieve a solid improvement (~0.6 BLEU points) over the baseline despite including weak ensemble components (e.g., Finnish pivoting, which is almost 5 BLEU points lower than direct translation).
-If we limit the ensemble components to the strong ones (top-3), EBBS can achieve an even higher improvement of ~0.8 BLEU points.
-This analysis highlights the robustness of our EBBS algorithm.
+We also evaluate EBBS-based distillation (Row 8, Table 1).
+Again, since Europarl has 56 zero-shot directions, we follow the standard practice (Fan et al.
+2021) and select a subset of directions, namely, Danish to other languages, to save computational cost.
+As seen in Row 8, EBBS-based distillation consistently achieves the highest performance in all directions (except for Danish-to-Dutch translation).
+This shows that an EBBS-distilled model can outperform EBBS, which is not surprising because learning can smooth out the noise of various heuristics (Deshmukh et al.
+2021; Jolly et al.
+2022), such as the ensemble algorithm in our scenario.
+Importantly, EBBS-based distillation achieves significantly higher translation quality with no inference overhead compared with direct translation.
+Distillation analysis.
+We compare EBBS-based distillation with other distilling methods.
+Here, we only focus on Italian-to-Dutch [^4] We could only afford one translation direction for this analysis, because we need to train different models for all competing distilling methods.
+This differs from Table 1, where we follow previous work and perform EBBS-based distillation for Danish to other languages.
+We chose Italian-to-Dutch translation here, because it is the first one in IWSLT, conveniently also available in Europarl.
+translation to save computational cost.
+In particular, we consider two alternative distilling methods: direct and union distillation.
+Direct distillation finetunes the multilingual model with its own predictions based on direct translation.
+Union distillation, on the other hand, takes the union of the teachers’ outputs (direct and pivot translations) for training, which is under a controlled experimental setup, because it uses exactly the same translation paths as our EBBS-based distillation.
+As seen in Table 3, both direct and union distillation marginally improve the performance compared with no distillation.
+Intriguingly, learning from the union of multiple teachers is not necessarily better than learning from the best teacher (namely, direct translation).
+This is because multiple teachers may provide conflicting training signals and confuse the student model.
+On the contrary, our EBBS-based distillation consistently outperforms direct and union distillation on both datasets.
+This shows that our ensemble-then-distill approach is able to consolidate the knowledge of multiple teachers to better train the student model.
+Dataset Method BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
+IWSLT EBBS 19.52 51.87 25.12 13.88 8.02 45.63 71.36 0.7341
+Direct Translation No distillation 17.46 50.49 23.01 12.01 6.66 43.73 72.02 0.7088
+Direct distillation 18.10 50.37 23.53 12.63 7.17 44.48 72.86 0.7144
+Union distillation 17.80 49.21 23.01 12.51 7.10 44.93 75.92 0.7221
+EBBS distillation 20.13 53.20 26.06 14.33 8.26 46.46 69.28 0.7428
+Europarl EBBS 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
+Direct Translation No distillation 25.33 56.32 30.08 19.01 12.78 52.32 66.56 0.8276
+Direct distillation 25.44 56.54 30.28 19.13 12.79 52.61 66.34 0.8286
+Union distillation 25.53 56.58 30.34 19.18 12.91 52.63 66.27 0.8282
+EBBS distillation 25.92 56.76 30.68 19.57 13.24 52.73 66.04 0.8307
+Table 3: Comparison of various distilling methods for Italian-to-Dutch translation. ↑/↓The higher/lower, the better.
+Method BLEU chrF2++ TER COMET
+Direct translation 25.33 52.32 66.56 0.8276
+Pivoting (en) 25.08 51.92 66.24 0.8322
+Pivoting (es) 24.40 51.71 67.91 0.8192
+Pivoting (pt) 24.34 51.61 67.68 0.8191
+Pivoting (fr) 24.20 51.61 67.84 0.8208
+Pivoting (de) 23.65 50.70 67.89 0.8157
+Pivoting (da) 23.12 50.36 69.00 0.8156
+Pivoting (fi) 20.74 48.11 70.59 0.8051
+Our EBBS 26.10 52.75 65.63 0.8340
+Table 4: The performance of direct/pivot translation and our EBBS for Italian-to-Dutch translation on Europarl.
+Further, the analysis suggests that our EBBS-distilled model achieves a speedup of multiple times compared with EBBS, because after distillation the model is used by direct translation.
+This is a significant result, because our EBBS-based distillation not only speeds up the EBBS ensemble approach, but also improves the translation quality of EBBS as shown in Row 8, Table 1.
+Analysis of ensemble components.
+In Table 4, we analyze the ensemble components to better understand our ensemble technique for zero-shot machine translation.
+As seen, direct translation is an effective approach, which is consistent with previous literature (Fan et al.
+2021; Liu et al.
+2021).
+English pivoting achieves higher performance for some metrics but lower for others; it is not conclusively better than direct translation, probably because of the error accumulation problem.
+Pivoting through non-English languages degrades the performance to a large extent because lacking supervision along the pivoting path leads to two steps of zero-shot translation.
+EBBS, on the other hand, combines the strengths of individual components and consistently outperforms them in all metrics.
 Figure 2: Analysis of the number of ensemble components for Italian-to-Dutch translation on Europarl.
@@
 Figure 2 shows the trends of incremental ensembles.
-If we add best pivot directions, the performance peaks at three ensemble components; interestingly, the inclusion of weaker components does not affect EBBS much.
-On the other hand, adding the worst pivot translation leads to an immediate drop of 1.6 BLEU points, which then largely recovers with the second pivot.
+If we add the best pivot directions, the performance peaks at three ensemble components; interestingly, the inclusion of weaker components does not affect EBBS much.
+On the other hand, adding the worst pivot translation at the beginning leads to an immediate drop of 1.6 BLEU points, which then largely recovers with the second pivot.
 This is reasonable because the worst pivot (Finnish) is 4.6 BLEU points lower than direct translation, and EBBS cannot decide on which of the two ensemble components to trust; despite this, the performance of EBBS is still much better than the average performance of the components.
-With a second pivot, it gives a third “opinion” when the first two components “disagree.” The performance continues to rise if more and stronger components are added.
+With a second pivot, there is a third “opinion” when the first two components “disagree.” The performance continues to rise if more and stronger components are added.
 In fact, our ensemble even surpasses the baseline with 4 weakest pivot translations, each of which is at least 1 BLEU point lower than the baseline.
 This demonstrates that EBBS is flexible and works well with both strong and weak ensemble components.
-Distillation results.
-We apply distillation techniques to our ensemble outputs to speed up inference.
-We follow the same ensemble setting (three components) as in Table 1, but only focus on Italian-to-Dutch translation because generating the training samples for all 66 zero-shot directions (each containing more than 100K samples) requires significant computing resources.
-We compare our EBBS-based distillation with two alternatives.
-A straightforward way to distill from multiple teachers (here, translation paths) is to take the union of teachers’ outputs for training.
-We also include self-distillation, which does not utilize different translation paths but generates samples by standard beam search for self-training.
-As seen in Table 5, self-distillation improves zero-shot translation on both IWSLT and Europarl.
-This is because the base model has never learned from parallel data between Italian and Dutch, and the self-distillation process enables the multilingual model to further refine its translation capabilities in this zero-shot direction.
-Similarly, we observe that union distillation also improves performance to some degree, but it is not better than self-distillation.
-This is consistent with the challenges of one-to-many learning seen in previous work (Wen et al., 2023a, b).
-In our case, learning from the outputs of multiple teachers in a straightforward manner may confuse the student, failing to consolidate and transfer the teachers’ knowledge.
-In comparison, EBBS-based distillation consistently outperforms both self-distillation and union distillation, achieving near-EBBS performance on Europarl.
-For IWSLT, our model even outperforms EBBS, as IWSLT is a small dataset and the high-quality outputs generated by EBBS can largely improve the base model (no distillation) by 2.7 points.
-Method BLEU BLEU1 BLEU2 BLEU3 BLEU4
-IWSLT EBBS 19.52 51.87 25.12 13.88 8.02
-[…]
-[…]
-Union distillation 17.80 49.21 23.01 12.51 7.10
-EBBS distillation 20.13 53.20 26.06 14.33 8.26
-Europarl EBBS 26.10 57.07 31.00 19.76 13.28
-Direct No distillation 25.33 56.32 30.08 19.01 12.78
-Self-distillation 25.44 56.54 30.28 19.13 12.79
-[…]
-EBBS distillation 25.92 56.76 30.68 19.57 13.24
-Table 5: Comparison of various distillation methods for Italian-to-Dutch translation.
-We provide an efficiency analysis in Appendix D, showing that the inference time of EBBS is linear to the number of ensemble components.
-Overall, our EBBS-based distillation achieves a significant speed up while maintaining, or even improving, translation quality.
-Additional results. We show additional results in the appendices. C: Analysis of beam size, D: Analysis of inference efficiency, E: Entropy of distilled models, and F: Case study.
+Appendix. We present additional details and results in the appendix:
+  •  A.
+    Beam search,
+  •  B.
+    Experimental details,
+  •  C.
+    Analysis of inference efficiency,
+  •  D.
+    Average performance across tasks,
+  •  E.
+    Analysis of beam size,
+  •  F.
+    Entropy of distilled models,
+  •  G.
+    Analysis of voting methods in EBBS, and
+  •  H.
+    Case study.
 5 Conclusion
@@
 We evaluated EBBS on two popular zero-shot translation datasets, IWSLT and Europarl.
-Results show that EBBS outperforms existing ensemble techniques, and that the high-quality translations produced by EBBS can be used for distillation to improve translation efficiency (and sometimes also quality over EBBS).
-6 Limitations
-Our work features both algorithmic design and empirical effectiveness, but may also have limitations.
-First, our work focuses on zero-shot translation, due to the background of our project.
-Nevertheless, the EBBS algorithm that we have developed is a general decoding method and is potentially applicable to various text generation tasks.
-We are happy to explore this direction as future work, such as building ensembles of large language models.
-Second, our distillation experiments were conducted on Italian-to-Dutch translation because it appears in both IWSLT and Europarl datasets.
-We are unable to perform all-direction distillation in this project, because Europarl, say, has 56 zero-shot directions, and distillation techniques require extensive computing resources to generate pseudo-groundtruth sentences.
-Nevertheless, our distillation results are consistent on both datasets, and therefore, not including all-direction distillation does not hurt the validity of our method.
-Moreover, our work points to interesting future work, such as iteratively applying our ensemble-then-distill procedure to build a self-improving multilingual machine translation system.
-Third, we have mainly used BLEU (in particular, SacreBLEU) scores to evaluate translation quality, which is standard in multilingual translation research Johnson et al.
-(2017); Liu et al.
-(2021); Fan et al.
-(2021).
-We are aware of language model-based scores, such as BERTScore Zhang et al.
-(2019), for evaluating the translation in certain languages, but they may not be appropriate for low-resource languages and are not commonly used in our setting.
-We have not performed human evaluation either, due to both practical and ethical concerns.
-Practically, it is difficult to find qualified annotators because our multilingual setting requires an annotator to know a large number of languages in addition to English.
-Ethically, it is inappropriate to ask annotators to study these languages before or during annotation, which is exhausting and may last for extended periods of time.
-That being said, our work uses the most standard metrics and results are consistent in most BLEU variants on both datasets, with a strong statistical confidence level based on 62 translation directions.
-In addition, we have provided case studies to further illustrate how our EBBS works.
-Therefore, we deem our evaluation ethical, appropriate, and adequate.
-Acknowledgment
+Results show that EBBS outperforms existing ensemble techniques, and that the high-quality translations produced by EBBS can be used for distillation to improve translation efficiency (and sometimes also output quality).
+Acknowledgments
 The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), a Mitacs Accelerate project, the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, and the Digital Research Alliance of Canada (alliancecan.ca).
-In the title, we used the open-domain artwork “The Great Wave off Kanagawa” by Katsushika Hokusai.
+We used the open-domain artwork “The Great Wave off Kanagawa” by Katsushika Hokusai in the title and as the background for Figure 1.
 Appendix A Beam Search
@@
 Algorithm 2 Beam Search
-Appendix B Details of Our Experiments
-Dataset details. We evaluated our methods using IWSLT 2017 Cettolo et al. (2017) and Europarl v7 Koehn (2005). Table 6 provides a summary of the languages.
+Appendix B Experimental Details
+Dataset details. We evaluated our methods using IWSLT 2017 (Cettolo et al. 2017) and Europarl v7 (Koehn 2005). Table 5 provides a summary of the languages.
 The IWSLT 2017 translation dataset features multilingual data derived from TED talks.
-We followed previous work and used a standard split for zero-shot evaluation (Dabre and Kurohashi, 2017; Liu et al., 2021).
+We followed previous work and used a standard split for zero-shot evaluation (Dabre and Kurohashi 2017; Liu et al.
+2021).
 In particular, IWSLT contains English-centric training data for Italian, Dutch, and Romanian, while evaluation is performed in six zero-shot directions.
@@
 ro Romanian $\checkmark$
-Table 6: Summary of the language codes.
+Table 5: The languages in the IWSLT and Europarl datasets.
 Europarl is a multilingual dataset crawled from the proceedings of the European Parliament.
-We again followed previous work (Liu et al., 2021) and evaluated our methods with a standard split for the zero-shot setting, containing English-centric data for eight languages with a total of 56 zero-shot evaluation directions.
-We specifically adopted the non-overlapping setting: in the original corpus, a sentence may be translated into multiple languages, and the non-overlapping setup chooses only one target translation for each input.
-This prevents potential data leaking problems.
+We again followed previous work (Liu et al.
+2021) and evaluated our methods with a standard split for the zero-shot setting, containing English-centric data for eight languages with a total of 56 zero-shot evaluation directions.
+We adopted their non-overlapping setting: in the original corpus, a sentence may be translated into multiple languages, and the non-overlapping setup chooses only one target translation for each input.
+This prevents potential data-leaking problems.
 Europarl contains more data and languages than IWSLT, which further tests our method’s ability to generalize across multiple languages.
@@
 In implementation, we used the sum of log-probabilities for numerical stability.
-Moreover, our EBBS is built on top of the popular fairseq framework (Ott et al., 2019), using their beam search implementation as the backbone.
+Moreover, our EBBS is built on top of the popular fairseq framework (Ott et al.
+2019), using their beam search implementation as the backbone.
 Consequently, we inherit standard beam search implementation techniques such as length normalization and max length constraints, which are not detailed in our pseudocode.
-Beam Size BLEU BLEU1 BLEU2 BLEU3 BLEU4
-1 24.98 56.75 30.08 18.84 12.46
-3 25.99 57.04 30.92 19.65 13.17
-[…]
-[…]
-[…]
+Figure 3: Inference time analysis on the test set of Italian-to-Dutch translation from Europarl.
+Experiments were conducted on an AMD EPYC 7313 CPU and an NVIDIA RTX A6000 GPU, with a batch size of 300 samples.
+Dataset Model BLEU chrF2++ TER COMET
+IWSLT Direct translation (our replication) 17.29 42.57 72.46 0.7242
+Pivoting (en) 16.19 42.79 80.39 0.7184
+Word-level averaging ensemble 17.28 42.63 72.66 0.7237
+Word-level voting ensemble 16.99 42.55 75.23 0.7162
+Sequence-level voting ensemble (MBR) 16.72 42.50 76.12 0.7170
+EBBS (ours) 18.24 43.66 73.50 0.7383
+Europarl Direct translation (our replication) 27.85 54.37 63.40 0.8426
+Pivoting (en) 27.75 54.08 63.03 0.8480
+Word-level averaging ensemble 27.78 54.45 63.80 0.8409
+Word-level voting ensemble 27.62 54.05 63.11 0.8409
+Sequence-level voting ensemble (MBR) 28.14 54.43 62.75 0.8452
+EBBS (ours) 28.44 54.64 62.48 0.8488
+Table 6: Average performance statistics across all zero-shot language pairs in IWSLT and Europarl.
+Lower Upper BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
+1 1 24.98 56.75 30.08 18.84 12.46 51.97 66.13 0.8255
+3 3 25.99 57.04 30.92 19.65 13.17 52.65 65.60 0.8326
+5 5 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
+7 7 26.10 57.06 30.98 19.74 13.30 52.75 65.69 0.8346
+9 9 26.12 57.02 30.99 19.78 13.31 52.79 65.76 0.8352
+5 1 25.06 56.70 30.10 18.88 12.49 52.07 66.20 0.8264
+5 2 25.63 56.90 30.56 19.29 12.87 52.48 65.92 0.8311
+5 3 25.94 56.93 30.87 19.61 13.14 52.69 65.79 0.8330
+5 4 26.03 57.02 30.95 19.69 13.21 52.74 65.75 0.8332
+5 5 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
 Table 7: Beam size analysis for Italian-to-Dutch translation on the Europarl dataset.
-Appendix C Analysis of Beam Size
-We analyze the effect of the beam size and show the results in Table 7.
-As seen, performance rises with a larger beam size and eventually plateaus at around five, which is consistent with the practice of standard beam search Meister et al.
-(2020).
-This analysis also shows the robustness of our EBBS algorithm.
-Based on this experiment and efficiency considerations, we used a beam size of five in our main experiments.
-It should be mentioned that in principle our lower-level and upper-level beams need not have the same size.
-In our experiments, we simply set them to be the same.
-We have not tried full-grid hyperparameter tuning because the focus of our paper is algorithmic design, and as shown, our proposed EBBS is mostly robust to the beam size.
-Figure 3: Inference time analysis on the test set of Italian to Dutch translation from the Europarl dataset.
-Experiments were conducted on an AMD EPYC 7313 CPU and an NVIDIA RTX A6000 GPU, with an inference batch size of 300 samples.
-Appendix D Analysis of Inference Efficiency
+Appendix C Analysis of Inference Efficiency
 We analyze the efficiency of our ensemble approach.
@@
 The trend shows that it is computationally feasible to build an ensemble of even more components.
-Further, the analysis suggests that our EBB-distilled models achieves a speedup of multiple times depending on the number of ensemble components.
-Recall from §4.3 that our our EBBS-based distillation is able to transfer the ensemble’s knowledge to a single multilingual model, without hurting, if not improving, the translation quality.
+Importantly, recall this linearly increasing inference time is mitigated by EBBS distillation (Row 8, Table 1), which is as fast as one component.
+Appendix D Average Performance across Tasks
+In addition to the main results (Table 1), we further show the average performance across the 6 and 56 translation directions on IWSLT and Europarl using various metrics.
+We omit EBBS distillation, because distilling across all directions requires significant computing resources.
+As shown in Table 6, EBBS consistently outperforms existing baselines, with the only exception for TER on the IWSLT dataset, where direct translation achieved a higher performance.
+This is potentially because the TER score is based on edit distance, which is not robust to word reordering.
+In general, results show EBBS consistently outperforms competing methods.
+Appendix E Analysis of Beam Size
+We analyze the effect of different beam sizes on our EBBS algorithm.
+First, we study the setting where the lower- and upper-level beam sizes are matched.
+As seen in the top half of Table 7, the performance tends to increase with a larger beam size and eventually plateaus at around five, which is consistent with the practice of standard beam search (Meister, Cotterell, and Vieira 2020).
+Further, we analyze the setting where the upper- and lower-level beam sizes are not matched.
+Generally, the upper-level beam size should not exceed the lower-level beam size, because otherwise the upper-level beam may not be fully filled by the ensemble components.
+As shown in the bottom half of Table 7, EBBS performs better with larger upper-level beam sizes.
+This is understandable because a larger upper-level beam allows EBBS to explore more candidates in general.
+Overall, our analysis shows that EBBS is robust and works well with a variety of beam sizes.
+Based on this experiment and efficiency considerations, we used a beam size of five for both upper- and lower-level beams in our main experiments.
 Dataset Method BLEU Entropy
-[…]
+IWSLT EBBS 19.52 -
 Direct translation No distillation 17.46 2.46
-Self-distillation 18.10 1.62
+Direct distillation 18.10 1.62
 Union distillation 17.80 1.80
@@
 Direct translation No distillation 25.33 2.06
-Self-distillation 25.44 1.44
+Direct distillation 25.44 1.44
 Union distillation 25.53 1.59
-[…]
+EBBS distillation 25.92 1.51
 Table 8: Entropy of various distillation techniques on Italian-to-Dutch translation.
+Voting scheme BLEU BLEU1 BLEU2 BLEU3 BLEU4 chrF2++ TER COMET
+None (beam search) 25.33 56.32 30.08 19.01 12.78 52.32 66.56 0.8276
+Total-sum 25.27 56.67 30.27 19.07 12.65 52.19 66.12 0.8311
+Max 25.81 56.89 30.76 19.51 13.09 52.46 65.92 0.8300
+0/1 25.84 56.99 30.78 19.49 13.05 52.61 65.80 0.8322
+Top-$Z$ sum (ours) 26.10 57.07 31.00 19.76 13.28 52.75 65.63 0.8340
+Table 9: Comparison of different ensemble variants, using Italian-to-Dutch translation in the Europarl dataset as the testbed.
+Appendix F Entropy of Distilled Models
+We would like to understand why EBBS-based distillation largely outperforms other methods, such as union distillation (§4.3).
+Our hypothesis is that cross-entropy distillation loss with diverse samples may lead to an overly smooth distribution, which in turn would affect the model performance (Wen et al.
+2023b; Shayegh et al.
+2024).
+We show the average prediction entropy of our distilled models in Table 8. For some input $\mathbf{x}$ and generation step $t$, the prediction entropy is
+$H=-\sum_{\mathrm{y}\in V}$ $\;p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})\log p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})$
+A large entropy generally indicates that the model is less certain, producing a more uniform prediction, whereas a low entropy indicates that the model is confident, producing a more peaked distribution.
+As seen in Table 8, the model without distillation yields the highest entropy, suggesting that it is uncertain about zero-shot translation probably due to a lack of training signals.
+Union distillation trains the model from the union of ensemble components’ outputs.
+It reduces prediction entropy compared with no distillation, but due to the nature of cross-entropy loss, it remains the highest among all distillation variants.
+Direct distillation is based on direct translation only, reinforcing the model’s current belief and thus producing the lowest entropy.
+On the contrary, our EBBS-based distillation achieves a moderate entropy on both datasets.
+It should be emphasized that the entropy analysis merely shed light on how different distillation methods behave, but the entropy itself does not indicate the quality of a model.
+We quote BLEU scores from Table 3, which has suggested that our EBBS-based distillation achieves similar or higher performance compared with EBBS, consistently outperforming other distillation methods.
+Appendix G Analysis of Voting Methods in EBBS
+In our EBBS algorithm, the lower-level beams are synchronized into a shared upper-level beam by voting.
+Specifically, EBBS uses a mechanism of top-$Z$ sum voting, where we add the ensemble components’ probabilities for each appearance of a candidate in the lower-level beam, shown in Eqn.
+(3).
+Here, we analyze a few alternative voting methods for EBBS.
+If EBBS adopts total-sum voting, it still uses lower-level beams to find candidates, but adds all components’ probabilities together.
+This is equivalent to applying the common averaging ensemble to the top-$Z$ candidates.
+However, it differs from our approach, because in total-sum voting, a component will vote even if the candidate does not appear in its own lower-level beam; the probability after voting in Eqn.
+(3) is substituted with $\frac{1}{K}\sum_{k}p_{k}(\mathbf{y}|\mathbf{x})$.
+As shown in Table 9, EBBS with total-sum voting performs worse than direct translation, suggesting the importance of ignoring the components whose lower-level beam does not contain the candidate.
+This is analogous to nucleus sampling (Holtzman et al.
+2019), where the long tail of a distribution is mainly noise and should be ignored.
+Other voting schemes that EBBS may use include 0/1 voting and max voting.
+The former selects the candidates that appear most in the lower-level beams, disregarding the probability values (unless for ties); the latter chooses the maximum probability across the lower-level beams, which gives preference to sequences through a maximization bias (Hasselt 2010; van Hasselt, Guez, and Silver 2016).
+As seen, EBBS performs relatively well with both of these voting schemes, achieving a decent improvement over the baseline approach; however, their performance is worse than our top-$Z$ sum voting.
+Overall, the proposed bi-level beam search ensemble is effective with different voting schemes (except for the total-sum voting), and our top-$Z$ sum voting works the best among these variants.
 IWSLT
@@
 Table 10: Case studies, where the source language is Italian and the target is Dutch. We provide English interpretations in (italic) for non-English text using Google Translate.
-Appendix E Entropy of Distilled Models
-We would like to understand why EBBS-based distillation largely outperforms other methods, such as union distillation (§4.3).
-The hypothesis is that cross-entropy distillation loss with diverse samples may lead to an overly smooth distribution, which in turn would affect the model performance Wen et al.
-(2023b); Shayegh et al.
-(2024).
-We show the average prediction entropy of our distilled models in Table 8. For some input $\mathbf{x}$ and generation step $t$, the prediction entropy is
-$H=-\sum_{\mathrm{y}\in V}$ $\;p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})\log p(\mathrm{y}|\hat{\mathbf{y}}_{1:t-1},\mathbf{x})$
-Here, a large entropy generally indicates that the model is less certain, producing a more uniform prediction, whereas a low entropy indicates that the model is confident, producing a more peaked distribution.
-As seen in Table 8, the model without distillation yields the highest entropy, suggesting that it is uncertain about zero-shot translation probably due to a lack of training signals.
-Union distillation trains the model from the union of ensemble components’ outputs.
-It reduces prediction entropy compared with no distillation, but due to the nature of cross-entropy loss, it remains the highest among all distillation variants.
-Self-distillation is based on direction translation only, reinforcing the model’s current belief and thus producing the lowest entropy.
-On the contrary, our EBBS-based distillation achieves a moderate entropy on both datasets.
-It should be emphasized that the entropy analysis only merely light on how different distillation methods behave, but the entropy itself does not indicate the quality of a model.
-We quote BLEU scores from Table 5, which has suggested that our EBBS-based distillation achieves similar or higher performance compared with EBBS, consistently outperforming other distillation methods.
-Appendix F Case Study
+Appendix H Case Study
 Table 10 shows examples of direct, pivot, and EBBS translations.